### v1

# PF-LHM: 3D Animatable Avatar Reconstruction from Pose-free Articulated Human Images

Abstract
Reconstructing an animatable 3D human from casually captured images of an articulated subject without camera or human pose information is a practical yet challenging task due to view misalignment, occlusions, and the absence of structural priors. While optimization-based methods can produce high-fidelity results from monocular or multi-view videos, they require accurate pose estimation and slow iterative optimization, limiting scalability in unconstrained scenarios. Recent feed-forward approaches enable efficient single-image reconstruction but struggle to effectively leverage multiple input images to reduce ambiguity and improve reconstruction accuracy. To address these challenges, we propose PF-LHM, a large human reconstruction model that generates high-quality 3D avatars in seconds from one or multiple casually captured pose-free images. Our approach introduces an efficient Encoder-Decoder Point-Image Transformer architecture, which fuses hierarchical geometric point features and multi-view image features through multimodal attention. The fused features are decoded to recover detailed geometry and appearance, represented using 3D Gaussian splats. Extensive experiments on both real and synthetic datasets demonstrate that our method unifies single- and multi-image 3D human reconstruction, achieving high-fidelity and animatable 3D human avatars without requiring camera and human pose annotations. Code and models will be released to the public.
Figure 1: 3D Avatar Reconstruction and Animation Results of our PF-LHM. Given a set of $N\geq 1$ images of a human subject, without requiring camera parameters or human pose annotations, our method can reconstruct a high-fidelity, animatable 3D human avatar in seconds.
1 Introduction
Reconstructing high-quality, animatable 3D human avatars from casually captured images is a crucial task in computer graphics, with broad applications like virtual reality and telepresence. A practical solution should support rapid and robust reconstruction from minimal input—ideally using only one or a few casually captured images, without relying on camera parameters, human pose annotations, or controlled capture environments. Such capability is essential for enabling scalable and accessible avatar generation in real-world scenarios.
Existing approaches to animatable 3D human reconstruction from monocular or multi-view videos typically rely on optimization-based frameworks that minimize photometric or silhouette reprojection losses [2, 45, 8]. These methods usually require dozens or hundreds of images with accurate human pose estimation as a prerequisite. Moreover, the optimization process is often computationally expensive, taking several minutes or even hours to converge, thus limiting real-time applications.
More recently, LHM [31], a feed-forward network for single-image 3D human reconstruction, has shown promising progress toward real-time performance. It employs a transformer-based architecture to fuse geometric point features initialized from the canonical SMPL-X surfaces and image features to directly predict a 3D Gaussian Splatting [17] based avatar from a single image. However, as single-image-based methods are inherently limited by partial observations, they often struggle to reconstruct occluded or unseen regions, leading to oversmoothed surfaces or noticeable artifacts [34, 60].
A straightforward extension of LHM [31] to multi-image settings would involve concatenating image tokens from multiple images and performing attention fusion. However, such a naive approach suffers from substantial memory and computational overhead due to the large number of geometric point features and the quadratic complexity of dense self-attention mechanisms.
In this work, we propose PF-LHM, a novel feed-forward framework for fast and high-fidelity 3D human reconstruction from one or a few images without requiring the camera and human poses. To achieve this, we design an efficient Encoder-Decoder Point-Image Transformer (PIT) Framework that hierarchically fuses 3D geometric features with multi-image cues. The framework is built upon Point-Image Transformer blocks (PIT-blocks), which enable interaction between geometric and image tokens via attention fusion while maintaining scalability through spatial hierarchy.
We start by representing the SMPL-X anchor points as geometric tokens and extracting image tokens from each input image. The encoder stage comprises several PIT-blocks to progressively downsample the geometric tokens via Grid Pooling [47]. At each layer, the downsampled point tokens interact with image tokens through multimodal attention [6], allowing compact yet expressive geometric representations to be enriched with visual information from multiple images. The decoder stage upsamples the geometric tokens to recover spatial resolution.
The resulting 3D geometry tokens are decoded to predict Gaussian splatting parameters, enabling photorealistic rendering and animation. To enhance robustness and generalization, we train our model on large-scale real-world human video datasets covering diverse clothing styles, body shapes, and viewing conditions.
In summary, our contributions are:
  • We introduce PF-LHM, to the best of our knowledge, the first feed-forward model capable of reconstructing high-quality, animatable 3D human avatars in seconds from one or a few casually captured images, without requiring either camera poses or human pose annotations.
  • We propose a novel Encoder-Decoder Point-Image Transformer (PIT) architecture that hierarchically fuses 3D geometric point features and 2D image features using multimodal attention, enabling efficient and scalable integration of multi-image cues.
  • Extensive experiments on both synthetic and real-world data demonstrate that PF-LHM unifies single- and multi-image 3D human reconstruction, with superior generalization and visual quality.
2 Related Work
2.1 Human Reconstruction from A Single Image
For single-image 3D human reconstruction, many methods adopt implicit neural representations [33, 34, 50, 4, 59, 57, 49, 53] to model complex human geometries. To improve geometric consistency and generalizability, some approaches [5, 16, 1, 3] rely on parametric body models such as SMPL [21, 26] to predict geometric offsets for the reconstruction of clothed humans. However, reconstruction from a single image is an ill-posed problem. Current cascade-type approaches [19, 44, 46, 39, 32] attempt to mitigate this issue by decoupling the process into two stages: multi-view image synthesis using generative models, followed by 3D reconstruction. While these methods require view-consistent generation in the first stage, which is often unstable and challenging, this ultimately affects the quality of the reconstruction.
Inspired by the success of large reconstruction models [10, 37], emerging solutions aim to enable direct generalizable reconstruction through feed-forward networks which significantly accelerate the inference time. Human-LRM [46] employs a feed-forward model to decode the triplane NeRF representation, then followed by a conditional diffusion-based novel views generation and reconstruction. IDOL [60] introduces a UV-Alignment transformer model to decode Gaussian attribute maps in a structured 2D UV space. LHM [31] leverages a Body-Head multimodal transformer architecture produces animatable 3D avatars with the face identity preservation and fine detail recovery. While these single-view methods often face challenges with occlusions and invisible regions, frequently resulting in geometrically implausible results or blurred textures, the proposed PF-LHM leverages a variable number of pose-free images to reconstruct photorealistic and animatable avatars.
Concurrent work, GIGA [61], introduces a generalizable human reconstruction model based on UV map representations. However, in contrast to our approach, GIGA requires sparse-view inputs where the same action is captured from multiple viewpoints, along with complex camera setups and motion calibration. These constraints make it challenging to apply in causal scenarios.
2.2 Human Reconstruction from Monocular Videos
Video-based techniques further improve reconstruction consistency by using temporal cues. 4D replay methods [25, 45] can reconstruct dynamic humans from monocular video or multiview video sequences, however, they are not able to drive the humans in novel poses since they do not build a standalone 3D model for humans. Therefore, a series of monocular video-based methods [13, 30, 12, 36] build a static 3D human model and can drive the human in novel poses by binding the skinning weight.
Another series of works [14, 23, 11, 8, 15, 29, 55] take it further by incorporating a 3D parametric human model into the optimization process, and thus can drive the human reconstruction in novel poses without any post-processing. Despite impressive visual fidelity, they often require dozens of minutes and dozens of views for a good optimization, which limits their practical usage in real-world scenarios.
Unconstrained collection is ideal input for a practical application. However, existing methods [51, 54] share a similar pipeline that uses a view generative model and score distillation sampling [27] for shape optimization. As a result, they are costly for offline training and impractical for online reconstruction. Paving in a new way, PF-LHM infers a human avatar from one, a few, or dozens of views under any poses in a feed-forward manner and costs only seconds, which is extremely efficient for online applications. Moreover, PF-LHM greatly outperforms any previous state-of-the-art methods on 3D human reconstruction and offers a more flexible input manner for the community.
2.3 Feed-Forward Scene Reconstruction
Recent years have witnessed a paradigm shift in geometric 3D vision, driven by the emergence of methods that eliminate traditional dependencies on camera calibration and multi-stage pipelines. At the forefront of this revolution lies the DUSt3R [42] framework, which reimagines 3D reconstruction as a direct regression problem from image pairs to 3D pointmaps. By discarding the need for intrinsic camera parameters, extrinsic pose estimation, or even known correspondence relationships, DUSt3R and its successors [52, 38, 22, 40] have democratized 3D vision, enabling rapid reconstruction across diverse scenarios while achieving state-of-the-art performance in depth estimation, relative pose recovery, and scene understanding. However, general feed-forward reconstruction methods assume that images are captured from a static scene [41, 18], while our PF-LHM can accept multiple human images with different camera and human poses as input and produce an animatable 3D avatar.
Table 1: Comparison with state-of-the-art 3D human reconstruction methods. FF stands for Feed-forward, PF for Pose-free, and AM for Animatable.
Method # Image FF PF AM Runtime
IDOL [60] $1$ Seconds
AniGS [32] $1$ 15 Minutes
LHM [31] $1$ Seconds
Vid2Avatar [7] $>100$ 1-2 Days
GPS-Gaussian [58] $2$ Seconds
3DGS-Avatar [28] $>20$ 0.5 Hours
InstantAvatar [14] $>20$ Minutes
GaussianAvatar [11] $>20$ 0.5-6 Hours
ExAvatar [23] $>20$ 1.5-5 Hours
PuzzleAvatar [51] $4\sim 6$ 4-6 Hours
Vid2Avatar-Pro [8] $>100$ Hours
Ours $\geq 1$ Seconds
Figure 2: Overview of the proposed PF-LHM. In the 2D space, we extract image tokens $\mathbf{T}_{\text{Img}}$ by DINOv2 from the input RGB images without camera parameters or human poses, which are then concatenated with deformation tokens $\mathbf{T}_{\text{Def}}$ to form 2D tokens $\mathbf{T}_{\text{2D}}$. In the 3D space, geometric tokens $\mathbf{T}_{\text{3D}}$ are represented by the MLP output of SMPL-X anchor points. Subsequently, we build our Encoder-Decoder Point-Image Transformer (PIT) to hierarchically fuse 3D tokens with 2D tokens, where the downsampled 3D tokens interact with 2D tokens via multimodal attention in each layer. The finalized 3D tokens are decoded to directly predict 3D Gaussian parameters, enabling animation and photorealistic rendering.
3 Method
3.1 Overview
Problem Formulation
Given a set of $N\geq 1$ RGB images of a human subject, without known camera parameters or human pose annotations, our goal is to reconstruct a high-fidelity, animatable 3D human avatar in seconds.
We adopt the 3D Gaussians splatting (3DGS) [17] as the representation, which allows for photorealistic, real-time rendering and efficient pose control. Each 3D Gaussian primitive is parameterized by its center location $\mathbf{p}\in\mathbb{R}^{3}$, directional scales $\bm{\sigma}\in\mathbb{R}^{3}$, and orientation (represented as a quaternion) $\mathbf{r}\in\mathbb{R}^{4}$. In addition, the primitive includes opacity $\rho\in[0,1]$ and spherical harmonic (SH) coefficients $\mathbf{f}$ to model view-dependent appearance.
Inspired by LHM [31], we employ a set of spatial points $P\in\mathbb{R}^{N_{\text{points}}\times 3}$ uniformly sampled from the SMPL-X surface in its canonical pose to serve as the anchors. Conditioned on the multi-image inputs, these points are processed and decoded to regress the human 3D Gaussian appearance in canonical space through a feed-forward transformer-based architecture. The pipeline can be formulated as:
$\mathbf{\chi}\{\mathbf{p},\mathbf{r},\mathbf{f},\rho,\sigma\}=\text{PF-LHM}(P\mid I^{1},\dots,I^{N}).$ (1)
Model Design
A straightforward solution to this problem is to extend LHM to support multiple image inputs by directly concatenating all available image tokens and performing attention operations between 3D point tokens and image tokens. However, this naive extension results in significant computational and memory overhead due to the quadratic complexity of self-attention operations with respect to the total number of tokens, i.e., $\mathcal{O}((N_{\textrm{points}}+N)^{2})$.
To mitigate this issue, we explore strategies to reduce the number of geometric point tokens involved in attention. However, we empirically observe that simply reducing the number of point tokens significantly degrades reconstruction performance. To address this trade-off, we propose an efficient Encoder-Decoder Point-Image Transformer Framework to fuse image features with geometric point features, as illustrated in Fig. 2, which maintains reconstruction quality while reducing the attention footprint.
The final geometric point features output from the decoder are utilized to regress 3D Gaussian parameters using lightweight multi-layer perceptron (MLP) heads. To account for non-rigid deformations such as clothing or hair, we introduce learnable deformation-aware tokens. These tokens, together with the geometric features, are used to predict residual offsets in canonical space. Finally, Linear Blend Skinning (LBS) is applied to animate the canonical avatar into the target pose.
3.2 Encoder-Decoder Point-Image Transformer Framework
To efficiently fuse multi-image features with 3D geometric information, we propose an encoder-decoder architecture based on Point-Image Transformer blocks (PIT-blocks). This framework enables hierarchical feature interaction while alleviating the computational and memory burden associated with dense attention.
We begin by projecting the SMPL-X anchor points in canonical space into a set of geometric tokens and encoding the input images into image tokens, as described in Sec. 3.3. The network consists of $N_{\text{layer}}$ PIT blocks, divided into an encoder stage and a decoder stage. In the first $\lfloor N_{\text{layer}}/2\rfloor$ encoder blocks, we progressively reduce the spatial resolution of the geometric tokens using Grid Pooling [47]. At each layer, the downsampled point tokens perform the attention operation with the image tokens, enabling compact geometric representations enriched with multi-image visual cues.
In the subsequent $\lceil N_{\text{layer}}/2\rceil$ decoder blocks, we upsample the geometric tokens to restore their original resolution. At each stage, the upsampled tokens are concatenated with the corresponding high-resolution features from the encoder via skip connections. These fused features are further refined by the PIT blocks to reconstruct detailed geometry and view-dependent appearance.
3.3 Geometric Point and Images Tokenization
Geometric Point Tokenization
To incorporate human body priors, we initialize a set of 3D query points $\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{points}}}\subset\mathbb{R}^{3}$ by uniformly sampling from the mesh of a canonical SMPL-X pose. Following the design of Point Transformer v3 (PTv3) [48], we first serialize these points into a structured sequence and then project them into a higher-dimensional feature space using an MLP. Formally, this process is expressed as:
$X$ $=\text{Serialization}(X),$ (2)
$\mathbf{T}_{\text{3D}}$ $=\text{MLP}_{\text{proj}}(X)\in\mathbb{R}^{N_{\text{points}}\times C_{\text{point}}},$
where $C_{\text{point}}$ denotes the dimensionality of the point tokens.
Multi-Image Tokenization
To obtain rich image features, we adopt DINOv2 [24], a vision transformer pretrained on large-scale in-the-wild datasets, as the image encoder $\mathcal{E}_{\text{Img}}$. Given an input image $I$, we extract a sequence of image tokens as follows:
$\mathbf{T}_{\text{Img}}=\mathcal{E}_{\text{Img}}(I)\in\mathbb{R}^{N_{\text{I}}\times C},$ (3)
where $N_{\text{I}}$ is the number of image tokens and $C$ is the output feature dimension of the transformer.
Deformation-aware Token Injection
To account for non-rigid deformations such as clothing and hair present in each image, we introduce a learnable deformation token specific to the observed subject, denoted as $\mathbf{T}_{\text{Def}}\in\mathbb{R}^{1\times C}$. This token is concatenated with the image token sequence $\mathbf{T}_{\text{Img}}$, forming the multi-image tokens:
$\mathbf{T}_{\text{I}}=[\mathbf{T}_{\text{Img}};\mathbf{T}_{\text{Def}}]\in\mathbb{R}^{(N_{\text{I}}+1)\times C},$ (4)
where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes token-wise concatenation along the sequence dimension.
3.4 Point-Image Transformer Block
After obtaining both geometric and image tokens, we design an efficient Point-Image Transformer Block (PIT-block), which comprises three core attention modules to facilitate cross-modal interaction:
Point-wise Attention
To model self-attention among geometric tokens, we adopt the patch-based point transformer blocks from PTv3 [48]. This design enables cross-patch interactions via randomized shuffling of point orders, as detailed in the Supplementary Materials:
$\mathbf{T}_{\text{3D}}=\text{PTv3-Block}(\mathbf{T}_{\text{3D}}).$ (5)
Image-wise Attention
Given the image token sequence $\mathbf{T}_{\text{2D}}=\{\mathbf{T}^{1}_{\text{I}},\dots,\mathbf{T}^{N}_{\text{I}}\}\in\mathbb{R}^{N\times(N_{\text{I}}+1)\times C}$, we apply self-attention independently to the tokens of each image. This updates the features within each frame based on its own image tokens:
$\mathbf{T}_{\text{2D}}=\text{Self-Attention}(\mathbf{T}_{\text{2D}}).$ (6)
Point-Image Attention
After obtaining the updated features for both point-wise and frame-wise modalities, we develop a global point-image attention mechanism to fuse the point and multi-image tokens. Our model builds upon the powerful Multimodal-Transformer (MM-Transformer) [6] to efficiently merge features from different modalities.
To enhance global context representation in the input images, we utilize the class token $\mathbf{T}_{\text{cls}}$ extracted from the first frame as learnable global context features. Additionally, to align the dimensions of different modalities, we incorporate projection MLPs into both the input and output layers of the MM-Transformer (MM-T):
$\mathbf{\bar{T}}_{\text{2D}}$ $=\text{Flatten}(\mathbf{T}_{\text{2D}})\in\mathbb{R}^{N(N_{\text{I}}+1)\times C},$ (7)
$\mathbf{\bar{T}}_{\text{3D}}$ $=\text{MLP}_{\text{proj}}(\mathbf{T}_{\text{3D}})\in\mathbb{R}^{N_{\text{points}}\times C},$
[…]
$\mathbf{T}_{\text{3D}}$ $=\text{MLP}_{\text{uproj}}(\mathbf{\bar{T}}_{\text{3D}})\in\mathbb{R}^{N_{\text{points}}\times C_{\text{point}}},$
$\mathbf{T}_{\text{2D}}$ $=\text{UnFlatten}(\mathbf{\bar{T}}_{\text{2D}},N)\in\mathbb{R}^{N\times(N_{\text{I}}+1)\times C}.$
This global point-image attention module enables effective fusion of geometric and visual features by leveraging cross-modal attention.
3.5 3D Human Gaussian Parameter Prediction
Given the fused point tokens $\mathbf{T}_{\text{3D}}$ obtained from the encoder-decoder transformer framework, we predict the parameters of 3D Gaussians in the canonical human space using a lightweight MLP head:
$\{\Delta\mathbf{p}_{i},\mathbf{r}_{i},\mathbf{f}_{i},\rho_{i},\bm{\sigma}_{i}\}$ $=\text{MLP}_{\text{regress}}(\mathbf{T}_{\text{3D}}^{(i)}),$ (8)
$\mathbf{p}_{i}$ $=\mathbf{x}_{i}+\Delta\mathbf{p}_{i},\quad\forall i\in\{1,\dots,N_{\text{points}}\},$
where $\Delta\mathbf{p}_{i}\in\mathbb{R}^{3}$ denotes the predicted residual offset from the corresponding canonical SMPL-X vertex $\mathbf{x}_{i}$, and $\mathbf{r}_{i},\mathbf{f}_{i},\rho_{i},\bm{\sigma}_{i}$ are the Gaussian orientation, feature vector, opacity, and scale, respectively.
Pose Conditioned Deformation
Although the regressed canonical-space Gaussians can be animated to target poses using Linear Blend Skinning (LBS), modeling clothing deformations presents challenges due to their complex non-rigid motion patterns, which LBS is often unable to capture adequately. To overcome this limitation, we use a lightweight MLP to predict pose-dependent residual deformations.
Specifically, we derive a deformation-aware token $\mathbf{\bar{T}}_{\text{def}}$ by averaging the fused deformation-aware tokens $\mathbf{T}_{\text{def}}$ across all frames, and concatenate it with the SMPL parameters to modulate the geometric tokens using Adaptive Layer Normalization. These modulated features are then processed through a sequence of MLP layers to generate non-rigid residual deformations:
$\Delta\mathbf{p}_{i}^{\text{motion}}=\text{MLP}_{\text{motion}}\left(\text{AdaLN}(\mathbf{T}_{\text{points}}^{(i)},[\mathbf{\bar{T}}_{\text{def}};\bm{\theta}])\right),$ (9)
where $\bm{\theta}$ represents the SMPL pose parameters. The final posed positions are then obtained by adding both canonical offsets and motion-specific deformations before LBS is applied.
3.6 Loss Function
Our training strategy integrates photometric supervision from unconstrained video sequences with geometric regularization on Gaussian primitives. This hybrid optimization framework enables the learning of deformable human avatars without the need for explicit 3D ground-truth annotations.
To better capture complex clothing deformations, we adopt a diffused voxel skinning approach as proposed in [30, 20]. Given the predicted 3DGS parameters $\bm{\chi}=(\mathbf{p},\mathbf{r},\mathbf{f},\rho,\sigma)$, we transform the canonical avatar into target view space using voxel-based skinning.
Photometric Loss
We render the animated Gaussian primitives via differentiable splatting to obtain an RGB image $\hat{I}$ and an alpha mask $\hat{M}$, based on the target camera parameters. Supervision is applied through the following photometric loss:
$\mathcal{L}_{\text{photometric}}=\lambda_{\text{rgb}}\mathcal{L}_{\text{color}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{per}}\mathcal{L}_{\text{lpips}},$ (10)
where $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ are L1 losses on RGB and alpha values respectively, and $\mathcal{L}_{\text{lpips}}$ is a perceptual loss measuring high-frequency feature similarity. We set the corresponding weights as $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{mask}}=0.5$, and $\lambda_{\text{per}}=1.0$.
Gaussian Regularization Loss
Empirically, we observe that using only mask supervision tends to encourage overly large Gaussian scales, especially near object boundaries, which leads to blurred renderings. To counteract this issue, we propose a Mask Distribution Loss $\mathcal{L}_{\text{dis}}$, which encourages uniform Gaussian distributions within human masks and sharper boundary representation. This is achieved by rendering an auxiliary mask $M_{\text{dis}}$ with fixed Gaussian parameters (opacity $\rho=0.95$, scale $\sigma=0.002$), and applying a L1 loss between $M_{\text{dis}}$ and the ground-truth human mask.
Furthermore, to reduce ambiguities in canonical space supervision, we adopt two additional geometric regularizers from LHM [31]: (1) the As Spherical As Possible loss $\mathcal{L}_{\text{ASAP}}$, which promotes isotropy in the 3D Gaussians, and (2) the As Close As Possible loss $\mathcal{L}_{\text{ACAP}}$, which preserves spatial coherence among neighboring primitives.
The combined geometric regularization term is defined as:
$\mathcal{L}_{\text{reg}}=\lambda_{\text{dis}}\mathcal{L}_{\text{dis}}+\lambda_{\text{ASAP}}\mathcal{L}_{\text{ASAP}}+\lambda_{\text{ACAP}}\mathcal{L}_{\text{ACAP}},$ (11)
with empirically chosen weights: $\lambda_{\text{dis}}=0.5$, $\lambda_{\text{ASAP}}=20$, and $\lambda_{\text{ACAP}}=5$.
Overall Loss
The overall training objective combines photometric reconstruction accuracy with geometric regularization:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{photometric}}+\mathcal{L}_{\text{reg}}.$ (12)
4 Experiments
Table 2: Comparison experiments with sparse-view input methods on public benchmark.
views InstantAvatar [14] GaussianAvatar [11] ExAvatar [23] PF-LHM-L
PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time
2 19.324 0.929 0.136 87 s 13.572 0.863 0.180 90 s 25.534 0.953 0.049 8.5 m 26.544 0.962 0.026 1.65 s
4 20.860 0.928 0.129 143 s 18.201 0.916 0.112 134 s 26.563 0.956 0.043 15 m 26.979 0.962 0.025 2.23 s
8 20.833 0.926 0.134 254 s 21.830 0.942 0.073 216 s 27.573 0.960 0.040 32 m 27.311 0.963 0.023 4.11 s
16 21.075 0.930 0.119 451 s 23.461 0.949 0.059 376 s 28.108 0.962 0.039 1.22 h 27.888 0.965 0.022 10.64 s
Table 3: Comparison experiments with sparse-view input methods on causal videos.
views InstantAvatar [14] GaussianAvatar [11] ExAvatar [23] PF-LHM-L
PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time
2 22.624 0.950 0.079 87 s 18.455 0.935 0.089 90 s 26.723 0.969 0.031 8.5 m 28.229 0.974 0.016 1.65 s
4 23.627 0.954 0.073 143 s 21.073 0.949 0.063 134 s 27.532 0.970 0.028 15 m 28.450 0.974 0.016 2.23 s
8 23.548 0.954 0.071 254 s 23.980 0.960 0.043 216 s 28.431 0.972 0.026 32 m 28.643 0.975 0.015 4.11 s
16 23.420 0.955 0.065 451 s 24.943 0.962 0.037 376 s 28.916 0.973 0.024 1.22 h 28.924 0.977 0.014 10.64 s
Figure 3: Qualitative results of animatable human reconstruction from sparse-view inputs.
Implementation Details
We design three variants of our model with $N_{\text{layer}}=4,6,8$ layers of the PI-MT block, corresponding to PF-LHM-S (small), PF-LHM-M (medium), and PF-LHM-L (large), respectively. The models contain approximately 500 MB, 700 MB, and 1000 MB training parameters in total. We train the model by minimizing the training loss using the AdamW optimizer for 60,000 iterations. A cosine learning rate scheduler is employed, with a peak learning rate of 0.0001 and a warm-up period of 3,000 iterations. During each batch, we randomly sample a number of frames in the range of $[1,16]$ from a randomly selected training video. Input images are resized to have a maximum dimension of 1024 pixels. Training is performed on 32 A100 GPUs over five days. To ensure training stability, we apply gradient norm clipping with a threshold of 0.1. Additionally, we utilize bfloat16 precision and gradient checkpointing to enhance GPU memory and computational efficiency.
Training Dataset
For our model training, we utilize approximately 300,000 in-the-wild video sequences collected from public video repositories, along with over 5,173 3D public synthetic static human scans sourced from 2K2K [9], Human4DiT [35], and RenderPeople. Specifically, we employ a sampling ratio of 19:1 to draw training batches from the in-the-wild and synthetic datasets to balance generalization and view-consistency. To address view bias in the video data, we sample from a diverse range of perspectives as uniformly as possible, guided by the estimated global orientation of SMPL-X.
Evaluation Protocol
We report PSNR, SSIM [43], and LPIPS [56] to assess rendering quality, and measure efficiency with GPU memory usage and inference time.
4.1 Comparison with Existing Methods
Animatable Human Reconstruction from Sparse Images
We conduct a comprehensive evaluation of PF-LHM by comparing it with three baseline methods for generating animatable human avatars from casually captured video sequences. We assess the efficiency and performance of our model using two types of datasets: one is a public benchmark that includes 20 video sequences from NeuMan [15], REC-MV [30], and Vid2Avatar [7], while the other comprises 24 casual video sequences collected via our smartphones.
Both Table 2 and Table 3 illustrate quantitative experiments evaluating our model against InstantAvatar [14], GaussianAvatar [11], and ExAvatar [23] on public and casual video sequences. Compared to the state-of-the-art (SOTA) baseline ExAvatar, our approach not only significantly accelerates the inference time but also yields comparable quantitative results. Specifically, for the model’s efficiency, our model promptly creates animable avatars in seconds while ExAvatar requires approximately 15 minutes to 1.2 hours, depending on the number of input images. In terms of model performance, SOTA methods typically require dozens of input images to achieve satisfactory metrics; Moreover, our model achieves more accurate results with substantially fewer inputs, and this capability improves as the number of input images increases. As shown in Figure 3, sparse input views lead to noticeable reconstruction artifacts with fitting-based frameworks, including geometric distortions and texture blurring. However, our PF-LHM achieves robust and high-fidelity reconstruction from sparse inputs and outperforms the SOTA baselines.
Table 4: Comparison with single-image method on pose animation using our in-the-wild dataset. * indicates we use 80000 query points in LHM for a fair comparison. + denotes the measured inference time of IDOL.
[…]
AniGS [32] 1 17.023 0.858 0.087 15 minutes 24 GB
IDOL [60] 1 17.835 0.871 0.083 1.93 seconds + 23 GB
LHM-0.7B* [31] 1 20.338 0.920 0.047 4.77 seconds 21 GB
PF-LHM-M 1 20.335 0.919 0.047 1.11 seconds 11 GB
2 20.891 0.923 0.042 1.42 seconds 11 GB
4 21.172 0.925 0.039 2.02 seconds 13 GB
8 21.574 0.926 0.037 3.7 seconds 15 GB
16 21.752 0.928 0.037 9.15 seconds 18 GB
Animatable Human Reconstruction from a Single Image
We evaluate PF-LHM against three baseline approaches for single-view animatable human reconstruction. The first baseline is AniGS [32], which employs a multi-view diffusion model to create canonical human avatars, followed by 4D Gaussian splatting (4DGS) optimization to address inconsistencies across different views. The second one is IDOL [60], using a UV-based transformer model to create the animatable avatars training from synthetic human datasets built on a human video diffusion model. The last baseline, LHM [31], introduces novel body-head transformer blocks that directly regress human Gaussian parameters in canonical space. For our evaluation, we employ 400 in-the-wild video sequences featuring individuals of various age groups, including young men and women, older adults, and children. For a fair comparison, we compare LHM with the same parameters and query points. Figure 7 shows a qualitative comparison experiments with LHM on in-the-wild data. The figure indicates that PF-LHM achieves results comparable to LHM with single-view image input. Furthermore, as the number of inputs increases, our method generates increasingly realistic and detailed results.
Table 4 presents two key findings regarding single-view human reconstruction. Our unified framework not only achieves competitive quantitative results compared to LHM but also the inference speed that is four times faster than that of LHM. Furthermore, as the number of input views increases, our method’s performance also improves. Specifically, in comparison to a single-image input, using 16 input views results in improvement of 1.417, 0.09, and 0.01 in PSNR, SSIM, and LPIPS, respectively. Figure 7 shows a qualitative comparison experiment with LHM.
4.2 Qualitative Results
As demonstrated in Fig. 8, we present the animation results of avatars generated from various in-the-wild monocular videos, including NeuMan [15], REC-MV [30], Vid2Avatar [7], MVHumanNet [49] and our causal video dataset. Our PF-LHM is capable of generalizing across different identities and garment styles, producing highly realistic renderings for novel human poses and arbitrary viewpoints.
Figure 4: Animatable human reconstruction comparisons from sparse images on in-the-wild videos.
PF-LHM-S PF-LHM-M PF-LHM-L GT
Figure 5: Ablation study on model design and parameters.
Figure 6: PSNR performance of our models with different input views.
Figure 7: Comparison results of animatable human reconstruction methods with LHM.
Figure 8: Visual animation results of avatars created from monocular in-the-wild videos. The created 3D avatars can be animated using novel human poses and demonstrate highly detailed appearance. Also, the figure effectively illustrates the generalization capability of our approach, delivering outstanding results across various demographics, including the elderly, children, men, and women, as well as individuals of differing heights and body types.
Table 5: Model Efficiency. We evaluate the training and inference efficiency of various backbones. The batch size is set to 1, the number of input views is equal to 16. ‘# Points’ refers to the number of geometric points, and ‘Time’ denotes the duration of a single iteration during both training and inference.
Methods Params. # Points Training Inference
Time Memory Time Memory
LHM-0.5 B [31] 500 MB 80 K 26.33 s 54.97 GB 21.61 s 27.8 GB
LHM-0.7 B [31] 700 MB 80 K 50.23 s 62.27 GB 43.06 s 28.3 GB
LHM-1.0 B [31] 1000 MB 80 K 70.61 s 71.55 GB 64.29 s 29.7 GB
LHM-1.0 B [31] 1000 MB 160 K - $>$ 80 GB 159.19 s 35.8 GB
PF-LHM-S 500 MB 80 K 8.77 s 46.5 GB 7.35 s 17.0 GB
PF-LHM-M 700 MB 80 K 10.01 s 50.4 GB 9.03 s 18.3 GB
PF-LHM-L 1000 MB 80 K 11.13 s 55.1 GB 10.11 s 19.2 GB
PF-LHM-L 1000 MB 160 K 11.20 s 55.8 GB 10.64 s 20.0 GB
4.3 Ablation Study
Model Efficiency
Table 5 presents quantitative results that substantiate the efficiency of our model, tested on NVIDIA A100-80G hardware. The table clearly illustrates that our novel framework significantly outperforms the original LHM architecture, achieving notably reduced training times and lower memory consumption. Furthermore, during testing, our inference times are about 5$\sim$10 times faster than those of LHM.
Model Parameter Scalability
To verify the scalability of our PF-LHM, we train variant models with increasing parameter numbers by scaling the layer numbers. Table 6 compares performance across various model capacities. Our experiments indicate that increasing the number of model parameters correlates with improved performance. Figure 6 presents the comparison among PF-LHM-S, PF-LHM-M, and PF-LHM-L where the larger model achieves more accurate reconstruction. Also, Fig. 6 depicts the performance of different models based on varying input view counts, clearly indicating that our model is scalable and that performance improves with an increased number of input images.
Number of Query Points
Table 6 shows an ablation study analyzing the effect of varying the number of query points on public video datasets. As the number of query points increases from 40K to 80K, our model demonstrates improvements in PSNR, SSIM, and LPIPS by 0.562, 0.006, and 0.002, respectively. However, when the number of query points is increased further from 80K to 160K, we observe a slight gain in performance. Therefore, we set the number of query points to 80K to achieve an optimal balance between efficiency and model performance.
Table 6: Analysis of model parameters and 3D geometric point numbers.
[…]
PF-LHM-M 40 K 28.001 0.969 0.017 7.32 seconds 17 GB
PF-LHM-M 160 K 28.565 0.975 0.014 10.64 seconds 20 GB
PF-LHM-S 80 K 28.089 0.970 0.017 7.35 seconds 17 GB
PF-LHM-M 80 K 28.563 0.975 0.015 9.03 seconds 18 GB
PF-LHM-L 80 K 28.924 0.977 0.014 10.11 seconds 19 GB
5 Conclusion
We present PF-LHM, a novel feed-forward framework for rapid and high-fidelity 3D human avatar reconstruction from one or a few casually captured, pose-free images. Our approach introduces the Encoder-Decoder Point-Image Transformer (PIT), which enables efficient multimodal fusion between geometric point tokens and sparse multi-view image tokens through hierarchical attention. By leveraging the PIT architecture, our method achieves superior scalability, generalization, and reconstruction quality while significantly reducing runtime compared to optimization-based baselines. Extensive experiments across synthetic and real-world datasets demonstrate that PF-LHM effectively unifies single- and sparse-view reconstruction and supports realistic avatar animation using 3D Gaussian Splatting.
Appendix A Implementation Details
A.1 Details of Points Serialization
To trade the scalability and efficiency of our feed-forward framework, we leverage serialization to transform unstructured SMPL-X anchor points into structured data format. Following Point Transformer v3 (PTv3) [48], we mixture 4 patterns of serialization, including Z-order, Hibert, Trans Z-order and Trans Hibert, and apply random shuffle to the order of serialization patterns before each PTv3 block.
Figure 9: Detailed architecture of PTv3 Block [48].
A.2 Details of PTv3 Block
As illustrated in Fig. 9, our point attention architecture integrates the mechanism introduced in Point Transformer v3 [48]. This architecture employs patch-based self-attention to accelerate the forward pass. Furthermore, Grip Pooling is utilized to downsample the point cloud, enhancing the efficiency of point cloud self-attention. Importantly, to facilitate interaction among different patch groups, a shuffling order is implemented.
In our setup, the patch sizes are set to 4096, 2048, and 1024 for the 80K point clouds, decreasing in accordance with the downsampling process. For the 160K point clouds, the patch sizes are configured to 8192, 4096, and 2048.
Table 7: Ablation study on dataset scalability. The metrics are calculated based on 16-view image inputs.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
PF-LHM-L + Synthetic Data 26.148 0.952 0.027
PF-LHM-L + 10K Videos 27.456 0.970 0.018
PF-LHM-L + 100K Videos 28.312 0.975 0.017
PF-LHM-L + All 28.924 0.977 0.014
Only Synthetic 10K Videos 100K Videos 300K Videos GT
Figure 10: Ablation study on dataset scalability.
Ground Truth w/o Deformation w/ Deformation
Figure 11: The ablation study for posed-conditioned deformation. The red boxes highlight the visual difference mentioned above.
Input w/o $\mathcal{L}_{\text{dis}}$ w/ $\mathcal{L}_{\text{dis}}$ Human Mask
Figure 12: The ablation study for mask distribution loss.
Appendix B Experiments
B.1 Details of Testing Dataset
For our qualitative results, we utilize videos from Vid2Avatar [7], specifically those indexed as: 00000_random, 00020_Dance, 00069_Dance, exstrimalik, and Yuliang. Additionally, we draw on videos from Rec-MV [30], including the following indices: anran_dance_self_rotated, anran_purple, anran_skirt, self-rotate-leyang, and xiaolin. Moreover, we use the NeuMan [15] dataset with the indices: citron and seattle. And, sparse view images are randomly sampled from MVHumanNet [49], with the following indices: 101330, 101425, 101483, 103383, 103528, 104013, 203243, and 103594, along with videos captured using our smartphone.
B.2 More Ablation Study
Details of Dataset Scalability
To assess the scalability of our dataset, we perform controlled experiments utilizing stratified random subsets of 10K and 100K from the original training dataset of 300K videos. Table 7 demonstrates that relying solely on the synthetic dataset leads to poor model generalization. In contrast, incorporating an in-the-wild dataset significantly improves the model’s robustness and performance during real-world evaluations. Furthermore, increasing the size of the dataset siginificantly improves model results, although the gains in performance tend to diminish with larger dataset sizes. Figure 10 illustrates the findings of our ablation study regarding dataset scalability.
Mask Distribution Loss
Gaussian models tend to learn large-scale Gaussian primitives rather than large step offsets. To address this issue, we introduce a mask distribution regularization loss that encourages Gaussian primitives to learn offsets instead of concentrating on large-scale axes. As illustrated in Fig. 12, employing this loss enables the mean position of the Gaussian primitives to be distributed more evenly across the ground truth mask area, thereby preventing the model from learning large-scale axes of the Gaussian primitives.
Posed-Condition Deformation
Figure 11 illustrates the impact of using pose-conditioned deformation. The results clearly indicate that our deformation tokens, which are conditioned on SMPL-X parameters, effectively model non-rigid deformations of loose garments, in contrast to the case without pose conditioning.
B.3 More Results
Figure 13 presents more animation results using eight input images. Our method enables high-fidelity reconstruction and animation of human avatars through the efficient Point-Image Transformer architecture, thereby demonstrating robust generalization capabilities and practical effectiveness.
Appendix C Limitations and Future Work
A primary limitation of our approach lies in its reliance on the SMPL-X template mesh for initializing geometric tokens. While this offers a strong structural prior, it can constrain reconstruction fidelity for subjects wearing loose or non-body-conforming garments, such as dresses, which deviate significantly from the SMPL-X topology. Additionally, due to the limited diversity of large motion poses in our training datasets, the model’s performance may degrade when encountering unseen or extreme poses. In future work, we plan to investigate more flexible garment representations and explore pose-independent anchor structures to better capture complex clothing dynamics and improve generalization to diverse motions.
Figure 13: More animation results of avatars created with 8-image inputs. Reference image is one of the input images.

### v2

@@
-PF-LHM: 3D Animatable Avatar Reconstruction from Pose-free Articulated Human Images
+LHM++: An Efficient Human Reconstruction Model for Pose-free Images to 3D
 Abstract
-Reconstructing an animatable 3D human from casually captured images of an articulated subject without camera or human pose information is a practical yet challenging task due to view misalignment, occlusions, and the absence of structural priors.
-While optimization-based methods can produce high-fidelity results from monocular or multi-view videos, they require accurate pose estimation and slow iterative optimization, limiting scalability in unconstrained scenarios.
-Recent feed-forward approaches enable efficient single-image reconstruction but struggle to effectively leverage multiple input images to reduce ambiguity and improve reconstruction accuracy.
-To address these challenges, we propose PF-LHM, a large human reconstruction model that generates high-quality 3D avatars in seconds from one or multiple casually captured pose-free images.
-Our approach introduces an efficient Encoder-Decoder Point-Image Transformer architecture, which fuses hierarchical geometric point features and multi-view image features through multimodal attention.
-The fused features are decoded to recover detailed geometry and appearance, represented using 3D Gaussian splats.
-Extensive experiments on both real and synthetic datasets demonstrate that our method unifies single- and multi-image 3D human reconstruction, achieving high-fidelity and animatable 3D human avatars without requiring camera and human pose annotations.
-Code and models will be released to the public.
-Figure 1: 3D Avatar Reconstruction and Animation Results of our PF-LHM.
+Reconstructing animatable 3D humans from casually captured images of articulated subjects without camera or pose information is highly practical but remains challenging due to view misalignment, occlusions, and the absence of structural priors.
+In this work, we present LHM++, an efficient large-scale human reconstruction model that generates high-quality, animatable 3D avatars within seconds from one or multiple pose-free images.
+At its core is an Encoder–Decoder Point–Image Transformer architecture that progressively encodes and decodes 3D geometric point features to improve efficiency, while fusing hierarchical 3D point features with image features through multimodal attention.
+The fused features are decoded into 3D Gaussian splats to recover detailed geometry and appearance.
+To further enhance visual fidelity, we introduce a lightweight 3D-aware neural animation renderer that refines the rendering quality of reconstructed avatars in real time.
+Extensive experiments show that our method produces high-fidelity, animatable 3D humans without requiring camera or pose annotations.
+Our code and project page are available at https://lingtengqiu.github.io/LHM++/.
+Figure 1: 3D Avatar Reconstruction and Animation Results of our LHM++.
 Given a set of $N\geq 1$ images of a human subject, without requiring camera parameters or human pose annotations, our method can reconstruct a high-fidelity, animatable 3D human avatar in seconds.
@@
 However, such a naive approach suffers from substantial memory and computational overhead due to the large number of geometric point features and the quadratic complexity of dense self-attention mechanisms.
-In this work, we propose PF-LHM, a novel feed-forward framework for fast and high-fidelity 3D human reconstruction from one or a few images without requiring the camera and human poses.
+In this work, we propose LHM++, a novel feed-forward framework for fast and high-fidelity 3D human reconstruction from one or a few images without requiring the camera and human poses.
 To achieve this, we design an efficient Encoder-Decoder Point-Image Transformer (PIT) Framework that hierarchically fuses 3D geometric features with multi-image cues.
@@
 The decoder stage upsamples the geometric tokens to recover spatial resolution.
-The resulting 3D geometry tokens are decoded to predict Gaussian splatting parameters, enabling photorealistic rendering and animation.
-To enhance robustness and generalization, we train our model on large-scale real-world human video datasets covering diverse clothing styles, body shapes, and viewing conditions.
+The resulting 3D geometry tokens are decoded to predict Gaussian splatting parameters, enabling high-quality rendering and animation.
+To further improve visual fidelity, we introduce a lightweight 3D-aware neural animation renderer, built on a DPT-head architecture, that refines the rendering quality of reconstructed avatars in real time.
+For robustness and generalization, we train our model on large-scale real-world human video datasets spanning diverse clothing styles, body shapes, and viewing conditions.
 In summary, our contributions are:
-  • We introduce PF-LHM, to the best of our knowledge, the first feed-forward model capable of reconstructing high-quality, animatable 3D human avatars in seconds from one or a few casually captured images, without requiring either camera poses or human pose annotations.
+  • We introduce LHM++, an efficient feed-forward model, capable of reconstructing high-quality and animatable 3D human avatars in seconds from one or a few casually captured images, without requiring either camera poses or human pose annotations.
   • We propose a novel Encoder-Decoder Point-Image Transformer (PIT) architecture that hierarchically fuses 3D geometric point features and 2D image features using multimodal attention, enabling efficient and scalable integration of multi-image cues.
-  • Extensive experiments on both synthetic and real-world data demonstrate that PF-LHM unifies single- and multi-image 3D human reconstruction, with superior generalization and visual quality.
+  • Extensive experiments on both synthetic and real-world data demonstrate that LHM++ unifies single- and multi-image 3D human reconstruction, with superior generalization and visual quality.
 2 Related Work
-2.1 Human Reconstruction from A Single Image
+Human Reconstruction from A Single Image
 For single-image 3D human reconstruction, many methods adopt implicit neural representations [40, 41, 61, 7, 73, 71, 60, 64] to model complex human geometries.
@@
 LHM [37] leverages a Body-Head multimodal transformer architecture produces animatable 3D avatars with the face identity preservation and fine detail recovery.
-While these single-view methods often face challenges with occlusions and invisible regions, frequently resulting in geometrically implausible results or blurred textures, the proposed PF-LHM leverages a variable number of pose-free images to reconstruct photorealistic and animatable avatars.
-Concurrent work, GIGA [61], introduces a generalizable human reconstruction model based on UV map representations.
-However, in contrast to our approach, GIGA requires sparse-view inputs where the same action is captured from multiple viewpoints, along with complex camera setups and motion calibration.
-These constraints make it challenging to apply in causal scenarios.
-2.2 Human Reconstruction from Monocular Videos
+While these single-view methods often face challenges with occlusions and invisible regions, frequently resulting in geometrically implausible results or blurred textures.
+Recently, HumanRAM [67] adapts human reconstruction to novel view and pose synthesis using LVSM with static pose sparse view inputs, achieving photorealistic results but suffering from slow transformer-based rendering.
+The concurrent work, GIGA [75], employs UV map representations for generalizable reconstruction, but it relies on multi-view captures of the same action, complex camera setups, and motion calibration.
+Human Reconstruction from Monocular Videos
 Video-based techniques further improve reconstruction consistency by using temporal cues.
@@
 Therefore, a series of monocular video-based methods [16, 36, 15, 45] build a static 3D human model and can drive the human in novel poses by binding the skinning weight.
-Another series of works [14, 23, 11, 8, 15, 29, 55] take it further by incorporating a 3D parametric human model into the optimization process, and thus can drive the human reconstruction in novel poses without any post-processing.
+Table 1: Comparison with state-of-the-art 3D human reconstruction methods. $F$ stands for Feed-forward, PF for Pose-free, and AM for Animatable.
+Method # Image $F$ PF AM Runtime
+CAR [25] 11 5 Minutes
+IDOL [74] 11 Seconds
+AniGS [38] 11 15 Minutes
+LHM [37] 11 Seconds
+Vid2Avatar [10] >100>100 1-2 Days
+Hugs [22] >80>80 30 Minutes
+Canonicalfusion [44] >1>1 11 Minutes
+GPS-Gaussian [72] 22 Seconds
+3DGS-Avatar [34] >20>20 0.5 Hours
+InstantAvatar [17] >20>20 Minutes
+GaussianAvatar [14] >20>20 0.5-6 Hours
+ExAvatar [29] >20>20 1.5-5 Hours
+PuzzleAvatar [62] $4\sim 6$ 4-6 Hours
+Vid2Avatar-Pro [11] >100>100 Hours
+Giga [75] $1\sim 4$ Seconds
+FRESA [52] $1\sim 4$ Seconds
+Ours $\geq 1$ Seconds
+Another series of works [17, 29, 14, 11, 18, 35, 66, 68, 65, 22] take it further by incorporating a 3D parametric human model into the optimization process, and thus can drive the human reconstruction in novel poses without any post-processing.
 Despite impressive visual fidelity, they often require dozens of minutes and dozens of views for a good optimization, which limits their practical usage in real-world scenarios.
@@
 As a result, they are costly for offline training and impractical for online reconstruction.
-Paving in a new way, PF-LHM infers a human avatar from one, a few, or dozens of views under any poses in a feed-forward manner and costs only seconds, which is extremely efficient for online applications.
-Moreover, PF-LHM greatly outperforms any previous state-of-the-art methods on 3D human reconstruction and offers a more flexible input manner for the community.
-2.3 Feed-Forward Scene Reconstruction
+Feed-Forward Scene Reconstruction
 Recent years have witnessed a paradigm shift in geometric 3D vision, driven by the emergence of methods that eliminate traditional dependencies on camera calibration and multi-stage pipelines.
@@
 By discarding the need for intrinsic camera parameters, extrinsic pose estimation, or even known correspondence relationships, DUSt3R and its successors [63, 47, 28, 50] have democratized 3D vision, enabling rapid reconstruction across diverse scenarios while achieving state-of-the-art performance in depth estimation, relative pose recovery, and scene understanding.
-However, general feed-forward reconstruction methods assume that images are captured from a static scene [41, 18], while our PF-LHM can accept multiple human images with different camera and human poses as input and produce an animatable 3D avatar.
-Table 1: Comparison with state-of-the-art 3D human reconstruction methods. FF stands for Feed-forward, PF for Pose-free, and AM for Animatable.
-Method # Image FF PF AM Runtime
-IDOL [60] $1$ Seconds
-AniGS [32] $1$ 15 Minutes
-LHM [31] $1$ Seconds
-Vid2Avatar [7] $>100$ 1-2 Days
-GPS-Gaussian [58] $2$ Seconds
-3DGS-Avatar [28] $>20$ 0.5 Hours
-InstantAvatar [14] $>20$ Minutes
-GaussianAvatar [11] $>20$ 0.5-6 Hours
-ExAvatar [23] $>20$ 1.5-5 Hours
-PuzzleAvatar [51] $4\sim 6$ 4-6 Hours
-Vid2Avatar-Pro [8] $>100$ Hours
-Ours $\geq 1$ Seconds
-Figure 2: Overview of the proposed PF-LHM.
-In the 2D space, we extract image tokens $\mathbf{T}_{\text{Img}}$ by DINOv2 from the input RGB images without camera parameters or human poses, which are then concatenated with deformation tokens $\mathbf{T}_{\text{Def}}$ to form 2D tokens $\mathbf{T}_{\text{2D}}$.
-In the 3D space, geometric tokens $\mathbf{T}_{\text{3D}}$ are represented by the MLP output of SMPL-X anchor points.
-Subsequently, we build our Encoder-Decoder Point-Image Transformer (PIT) to hierarchically fuse 3D tokens with 2D tokens, where the downsampled 3D tokens interact with 2D tokens via multimodal attention in each layer.
-The finalized 3D tokens are decoded to directly predict 3D Gaussian parameters, enabling animation and photorealistic rendering.
+However, general feed-forward reconstruction methods assume that images are captured from a static scene [51, 23], while our LHM++ can accept human images with different camera and human poses as input and produce an animatable 3D avatar.
+The concurrent work, FastVGGT [43], employs a token merge strategy to improve the efficiency of image token global attention.
+Figure 2: Overview of the proposed LHM++.
+In 2D space, we extract image tokens $\mathbf{T}_{\text{2D}}$ from input RGB images by DINOv2.
+In 3D space, geometric tokens $\mathbf{T}_{\text{3D}}$ are derived from SMPL-X anchor points via an MLP.
+Next, we design an Encoder-Decoder Point-Image Transformer (PIT) to hierarchically fuse 3D and 2D tokens, where the downsampled 3D tokens interact with 2D tokens via multi-modal attention in each layer.
+The final 3D tokens are decoded to predict 3D Gaussian parameters, followed by a light-weight DPT head for photorealistic animation.
 3 Method
@@
 Problem Formulation
-Given a set of $N\geq 1$ RGB images of a human subject, without known camera parameters or human pose annotations, our goal is to reconstruct a high-fidelity, animatable 3D human avatar in seconds.
+Given a set of $N\geq 1$ RGB images $I^{1},\dots,I^{N}$ of a human subject, without known camera parameters or human pose annotations, our goal is to reconstruct a high-fidelity, animatable 3D human avatar $\chi$ in seconds.
 We adopt the 3D Gaussians splatting (3DGS) [21] as the representation, which allows for photorealistic, real-time rendering and efficient pose control.
@@
 The pipeline can be formulated as:
-$\mathbf{\chi}\{\mathbf{p},\mathbf{r},\mathbf{f},\rho,\sigma\}=\text{PF-LHM}(P\mid I^{1},\dots,I^{N}).$ (1)
+$\mathbf{\chi}\{\mathbf{p},\mathbf{r},\mathbf{f},\rho,\bm{\sigma\}}=\text{LHM++}(P\mid I^{1},\dots,I^{N}).$ (1)
 Model Design
@@
 2, which maintains reconstruction quality while reducing the attention footprint.
+Additionally, with an increase in input images, image tokens dominate the attention computation.
+Inspired by ToME [5, 6], we merge the image tokens in a fixed ratio $r$ based on the key similarity map calculated from frame-wise attention before global attention, then unmerge them after global attention.
+Benefiting from these designs, our model significantly reduces the memory consumption and inference latency and achieves stable performance gains compared to a straightforward extension of LHM.
 The final geometric point features output from the decoder are utilized to regress 3D Gaussian parameters using lightweight multi-layer perceptron (MLP) heads.
-To account for non-rigid deformations such as clothing or hair, we introduce learnable deformation-aware tokens.
-These tokens, together with the geometric features, are used to predict residual offsets in canonical space.
-Finally, Linear Blend Skinning (LBS) is applied to animate the canonical avatar into the target pose.
+Subsequently, Linear Blend Skinning (LBS) is employed to animate the canonical avatar into the target pose.
+To enhance the quality of the animation rendering results, we develop a lightweight, real-time neural rendering network that improves rendering outcomes for novel views and poses based on 3D-aware human animation features.
 3.2 Encoder-Decoder Point-Image Transformer Framework
@@
 3.3.
-The network consists of $N_{\text{layer}}$ PIT blocks, divided into an encoder stage and a decoder stage.
+The encoder-decoder network consists of $N_{\text{layer}}$ PIT blocks.
 In the first $\lfloor N_{\text{layer}}/2\rfloor$ encoder blocks, we progressively reduce the spatial resolution of the geometric tokens using Grid Pooling [58].
@@
 Geometric Point Tokenization
-To incorporate human body priors, we initialize a set of 3D query points $\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{points}}}\subset\mathbb{R}^{3}$ by uniformly sampling from the mesh of a canonical SMPL-X pose.
+To incorporate human body priors, we initialize a set of 3D query points $\mathbf{X}=\left\{\mathbf{x}_{i}\right\}_{i=1}^{N_{\text{points}}}\subset\mathbb{R}^{3}$ by uniformly sampling from the mesh of a canonical SMPL-X pose.
 Following the design of Point Transformer v3 (PTv3) [59], we first serialize these points into a structured sequence and then project them into a higher-dimensional feature space using an MLP.
@@
 Given an input image $I$, we extract a sequence of image tokens as follows:
-$\mathbf{T}_{\text{Img}}=\mathcal{E}_{\text{Img}}(I)\in\mathbb{R}^{N_{\text{I}}\times C},$ (3)
+$\mathbf{T}_{\text{I}}=\mathcal{E}_{\text{Img}}(I)\in\mathbb{R}^{N_{\text{I}}\times C},$ (3)
 where $N_{\text{I}}$ is the number of image tokens and $C$ is the output feature dimension of the transformer.
-Deformation-aware Token Injection
-To account for non-rigid deformations such as clothing and hair present in each image, we introduce a learnable deformation token specific to the observed subject, denoted as $\mathbf{T}_{\text{Def}}\in\mathbb{R}^{1\times C}$.
-This token is concatenated with the image token sequence $\mathbf{T}_{\text{Img}}$, forming the multi-image tokens:
-$\mathbf{T}_{\text{I}}=[\mathbf{T}_{\text{Img}};\mathbf{T}_{\text{Def}}]\in\mathbb{R}^{(N_{\text{I}}+1)\times C},$ (4)
-where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes token-wise concatenation along the sequence dimension.
 3.4 Point-Image Transformer Block
@@
 Image-wise Attention
-Given the image token sequence $\mathbf{T}_{\text{2D}}=\{\mathbf{T}^{1}_{\text{I}},\dots,\mathbf{T}^{N}_{\text{I}}\}\in\mathbb{R}^{N\times(N_{\text{I}}+1)\times C}$, we apply self-attention independently to the tokens of each image.
+Given the image token sequence $\mathbf{T}_{\text{2D}}=\{\mathbf{T}^{1}_{\text{I}},\dots,\mathbf{T}^{N}_{\text{I}}\}\in\mathbb{R}^{N\times N_{\text{I}}\times C}$, we apply self-attention independently to the tokens of each image.
 This updates the features within each frame based on its own image tokens:
-$\mathbf{T}_{\text{2D}}=\text{Self-Attention}(\mathbf{T}_{\text{2D}}).$ (6)
+$\mathbf{T}_{\text{2D}},\mathbf{K}_{\text{2D}}=\text{Self-Attention}(\mathbf{T}_{\text{2D}}).$ (5)
+where $\mathbf{K}_{\text{2D}}\in\mathbb{R}^{NN_{\text{I}}\times C_{\text{attn}}}$ represents the key features with $C_{\text{attn}}$ channels from QKV in self-attention block, which are used to compute the similarity map of the image tokens.
 Point-Image Attention
@@
 Additionally, to align the dimensions of different modalities, we incorporate projection MLPs into both the input and output layers of the MM-Transformer (MM-T):
-$\mathbf{\bar{T}}_{\text{2D}}$ $=\text{Flatten}(\mathbf{T}_{\text{2D}})\in\mathbb{R}^{N(N_{\text{I}}+1)\times C},$ (7)
-$\mathbf{\bar{T}}_{\text{3D}}$ $=\text{MLP}_{\text{proj}}(\mathbf{T}_{\text{3D}})\in\mathbb{R}^{N_{\text{points}}\times C},$
-[…]
-$\mathbf{T}_{\text{3D}}$ $=\text{MLP}_{\text{uproj}}(\mathbf{\bar{T}}_{\text{3D}})\in\mathbb{R}^{N_{\text{points}}\times C_{\text{point}}},$
-$\mathbf{T}_{\text{2D}}$ $=\text{UnFlatten}(\mathbf{\bar{T}}_{\text{2D}},N)\in\mathbb{R}^{N\times(N_{\text{I}}+1)\times C}.$
-This global point-image attention module enables effective fusion of geometric and visual features by leveraging cross-modal attention.
+$\mathbf{\bar{T}}_{\text{3D}},\mathbf{\bar{T}}_{\text{2D}}$ $=\text{MLP}_{\text{proj}}(\mathbf{T}_{\text{3D}}),\text{Merge}(\text{Flatten}(\mathbf{T}_{\text{2D}}),\mathbf{K}_{\text{2D}}),$ (6)
+$\mathbf{\bar{T}}_{\text{3D}},\mathbf{\bar{T}}_{\text{2D}}$ $=\text{MM-T}(\mathbf{\bar{T}}_{\text{3D}},\mathbf{\bar{T}}_{\text{2D}}\mid\mathbf{T}_{\text{cls}}),$
+$\mathbf{T}_{\text{3D}},\mathbf{T}_{\text{2D}}$ $=\text{MLP}_{\text{uproj}}(\mathbf{\bar{T}}_{\text{3D}}),\text{UnFlatten}(\text{UnMerge}(\mathbf{\bar{T}}_{\text{2D}},\mathbf{K}_{\text{2D}}),N).$
+Specifically, $\text{MLP}_{\text{proj}}$ maps point tokens $\mathbf{T}_{\text{3D}}\in\mathbb{R}^{\bar{N}_{\text{points}}\times C_{\text{point}}}$ to $\mathbf{\bar{T}}_{\text{3D}}\in\mathbb{R}^{\bar{N}_{\text{points}}\times C}$, and $\text{MLP}_{\text{uproj}}$ performs the inverse mapping.
+The Merge and UnMerge operations are adapted from  [6], with a merge ratio $r=0.5$, reducing the number of image tokens from $NN_{\text{I}}$ to $NN_{\text{I}}/r$.
+This design reduces the time complexity of global attention from $\mathcal{O}((N_{\text{points}}+N)^{2})$ to $\mathcal{O}((\bar{N}_{\text{points}}+N/r)^{2})$.
 3.5 3D Human Gaussian Parameter Prediction
@@
 where $\Delta\mathbf{p}_{i}\in\mathbb{R}^{3}$ denotes the predicted residual offset from the corresponding canonical SMPL-X vertex $\mathbf{x}_{i}$, and $\mathbf{r}_{i},\mathbf{f}_{i},\rho_{i},\bm{\sigma}_{i}$ are the Gaussian orientation, feature vector, opacity, and scale, respectively.
-Pose Conditioned Deformation
-Although the regressed canonical-space Gaussians can be animated to target poses using Linear Blend Skinning (LBS), modeling clothing deformations presents challenges due to their complex non-rigid motion patterns, which LBS is often unable to capture adequately.
-To overcome this limitation, we use a lightweight MLP to predict pose-dependent residual deformations.
-Specifically, we derive a deformation-aware token $\mathbf{\bar{T}}_{\text{def}}$ by averaging the fused deformation-aware tokens $\mathbf{T}_{\text{def}}$ across all frames, and concatenate it with the SMPL parameters to modulate the geometric tokens using Adaptive Layer Normalization.
-These modulated features are then processed through a sequence of MLP layers to generate non-rigid residual deformations:
-$\Delta\mathbf{p}_{i}^{\text{motion}}=\text{MLP}_{\text{motion}}\left(\text{AdaLN}(\mathbf{T}_{\text{points}}^{(i)},[\mathbf{\bar{T}}_{\text{def}};\bm{\theta}])\right),$ (9)
-where $\bm{\theta}$ represents the SMPL pose parameters. The final posed positions are then obtained by adding both canonical offsets and motion-specific deformations before LBS is applied.
+3.6 3D-Aware Human Animation Rendering
+The final output of our model is to drive the reconstructed avatar into view space using the provided human motion control signal, specifically the SMPL-X parameters.
+Since our Gaussian model is deformed based on SMPL-X, it struggles to accurately represent loose clothing, such as skirts.
+To improve the rendering quality, we develop a lightweight, 3D-aware DPT head for generating final animation outcomes, inspired by LVSM [19] and HumanRAM [67].
+Unlike HumanRAM, which requires processing through transformers with multi-view images for each frame, making it computationally expensive.
+In contrast, our approach allows for real-time generation.
+The neural refinement process can be divided into two parts:
+Human Animation Feature Rasterization
+As we construct the canonical avatar in canonical space, we can leverage 3D priors to provide fine-grained control information.
+Specifically, given a reconstructed avatar obtained from Eq.
+7 and multi-scale geometric tokens $\mathbf{T}_{\text{3D}}$ from Eq.
+6, we employ Gaussian Rasterization [35], denoted as $\text{F}_{\text{GS}}$, to render high-dimensional geometric features into view space, resulting in $\mathbf{I}_{\text{feat}}\in\mathbb{R}^{H\times W\times C_{\textbf{point}}}$.
+Notably, as no 3D points are available to model the background, we utilize learnable background parameters $\mathbf{F}_{\text{back}}$, to represent background colors.
+The entire process can be formulated as follows:
+$\mathbf{I}^{l}_{\text{feat}}=\text{F}_{\text{GS}}(\mathbf{T}^{l}_{\text{3D}},\mathbf{r},\mathbf{p},\rho,\bm{\sigma};\mathbf{F}_{\text{back}}),$ (8)
+where $\mathbf{T}^{l}_{\text{3D}}$ represents the geometry features extracted from the $l$-th decoder layer of the PI Transformer framework.
+3D-Aware Neural Animation Rendering
+After obtaining the multi-scale 3D-aware human animation feature map $\mathbf{I}^{l}_{\text{feat}}$, we employ a lightweight neural network, as proposed in [39], to fuse these multi-scale animation features for novel view and pose rendering:
+$\mathbf{T}^{l}_{\text{feat}}$ $=\text{Patchify}(\mathbf{I}^{l}_{\text{feat}}),$ (9)
+$\hat{I}_{\text{NR}},\hat{M}_{\text{NR}}$ $=\text{DPT-head}(\mathbf{T}^{0}_{\text{feat}},\ldots,\mathbf{T}^{{N_{\text{layer}}/2}}_{\text{feat}}),$
+where Patchify refers to the operation that converts the 3D-aware animation features into image tokens using $8\times 8$ patch windows, $\hat{I}_{\text{NR}}$ and $\hat{M}_{\text{NR}}$ respectively representing the final predicted RGB image and mask results.
 3.7 Loss Function
@@
 To better capture complex clothing deformations, we adopt a diffused voxel skinning approach as proposed in [26, 36].
-Given the predicted 3DGS parameters $\bm{\chi}=(\mathbf{p},\mathbf{r},\mathbf{f},\rho,\sigma)$, we transform the canonical avatar into target view space using voxel-based skinning.
+Given the predicted 3DGS parameters, we transform the canonical avatar into target view space using voxel-based skinning.
 Photometric Loss
 We render the animated Gaussian primitives via differentiable splatting to obtain an RGB image $\hat{I}$ and an alpha mask $\hat{M}$, based on the target camera parameters.
+Additionally, we generate the neural rendering results $\hat{I}_{\text{NR}}$ and $\hat{M}_{\text{NR}}$ using a lightweight DPT head.
 Supervision is applied through the following photometric loss:
 $\mathcal{L}_{\text{photometric}}=\lambda_{\text{rgb}}\mathcal{L}_{\text{color}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{per}}\mathcal{L}_{\text{lpips}},$ (10)
-where $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ are L1 losses on RGB and alpha values respectively, and $\mathcal{L}_{\text{lpips}}$ is a perceptual loss measuring high-frequency feature similarity.
-We set the corresponding weights as $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{mask}}=0.5$, and $\lambda_{\text{per}}=1.0$.
+where $\mathcal{L}_{\text{color}}$ and $\mathcal{L}_{\text{mask}}$ represent L1 losses on the RGB and alpha values, corresponding to the 3D Gaussian splatting and neural rendering losses, respectively.
+The term $\mathcal{L}_{\text{lpips}}$ denotes a perceptual loss measures how similar the high-frequency features of neural-rendered and Gaussian Splat–rendered images are to those of the ground-truth images.
+The weights are set to $\lambda_{\text{rgb}}=1.0$, $\lambda_{\text{mask}}=0.5$, and $\lambda_{\text{per}}=1.0$.
 Gaussian Regularization Loss
-Empirically, we observe that using only mask supervision tends to encourage overly large Gaussian scales, especially near object boundaries, which leads to blurred renderings.
-To counteract this issue, we propose a Mask Distribution Loss $\mathcal{L}_{\text{dis}}$, which encourages uniform Gaussian distributions within human masks and sharper boundary representation.
-This is achieved by rendering an auxiliary mask $M_{\text{dis}}$ with fixed Gaussian parameters (opacity $\rho=0.95$, scale $\sigma=0.002$), and applying a L1 loss between $M_{\text{dis}}$ and the ground-truth human mask.
-Furthermore, to reduce ambiguities in canonical space supervision, we adopt two additional geometric regularizers from LHM [31]: (1) the As Spherical As Possible loss $\mathcal{L}_{\text{ASAP}}$, which promotes isotropy in the 3D Gaussians, and (2) the As Close As Possible loss $\mathcal{L}_{\text{ACAP}}$, which preserves spatial coherence among neighboring primitives.
-The combined geometric regularization term is defined as:
-$\mathcal{L}_{\text{reg}}=\lambda_{\text{dis}}\mathcal{L}_{\text{dis}}+\lambda_{\text{ASAP}}\mathcal{L}_{\text{ASAP}}+\lambda_{\text{ACAP}}\mathcal{L}_{\text{ACAP}},$ (11)
-with empirically chosen weights: $\lambda_{\text{dis}}=0.5$, $\lambda_{\text{ASAP}}=20$, and $\lambda_{\text{ACAP}}=5$.
+Although the photometric loss offers effective supervision in the target view space, the canonical representation remains under-constrained due to the ill-posed nature of monocular reconstruction.
+This limitation leads to deformation artifacts when warping the avatar into novel poses.
+To address this fundamental challenge, we employ Gaussian regularization loss $\mathcal{L}_{\text{reg}}$ to improve the learning of unobserved parts (see supplementary material for details).
 Overall Loss
@@
 4 Experiments
-Table 2: Comparison experiments with sparse-view input methods on public benchmark.
-views InstantAvatar [14] GaussianAvatar [11] ExAvatar [23] PF-LHM-L
+Table 2: Comparison with sparse-view input methods. The results are averaged on the public and smartphone-captured video sequences.
+views InstantAvatar [17] GaussianAvatar [14] ExAvatar [29] LHM++
 PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time
-2 19.324 0.929 0.136 87 s 13.572 0.863 0.180 90 s 25.534 0.953 0.049 8.5 m 26.544 0.962 0.026 1.65 s
-4 20.860 0.928 0.129 143 s 18.201 0.916 0.112 134 s 26.563 0.956 0.043 15 m 26.979 0.962 0.025 2.23 s
-8 20.833 0.926 0.134 254 s 21.830 0.942 0.073 216 s 27.573 0.960 0.040 32 m 27.311 0.963 0.023 4.11 s
-16 21.075 0.930 0.119 451 s 23.461 0.949 0.059 376 s 28.108 0.962 0.039 1.22 h 27.888 0.965 0.022 10.64 s
-Table 3: Comparison experiments with sparse-view input methods on causal videos.
-views InstantAvatar [14] GaussianAvatar [11] ExAvatar [23] PF-LHM-L
-PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time
-2 22.624 0.950 0.079 87 s 18.455 0.935 0.089 90 s 26.723 0.969 0.031 8.5 m 28.229 0.974 0.016 1.65 s
-4 23.627 0.954 0.073 143 s 21.073 0.949 0.063 134 s 27.532 0.970 0.028 15 m 28.450 0.974 0.016 2.23 s
-8 23.548 0.954 0.071 254 s 23.980 0.960 0.043 216 s 28.431 0.972 0.026 32 m 28.643 0.975 0.015 4.11 s
-16 23.420 0.955 0.065 451 s 24.943 0.962 0.037 376 s 28.916 0.973 0.024 1.22 h 28.924 0.977 0.014 10.64 s
-Figure 3: Qualitative results of animatable human reconstruction from sparse-view inputs.
-Implementation Details
-We design three variants of our model with $N_{\text{layer}}=4,6,8$ layers of the PI-MT block, corresponding to PF-LHM-S (small), PF-LHM-M (medium), and PF-LHM-L (large), respectively.
-The models contain approximately 500 MB, 700 MB, and 1000 MB training parameters in total.
-We train the model by minimizing the training loss using the AdamW optimizer for 60,000 iterations.
-A cosine learning rate scheduler is employed, with a peak learning rate of 0.0001 and a warm-up period of 3,000 iterations.
-During each batch, we randomly sample a number of frames in the range of $[1,16]$ from a randomly selected training video.
-Input images are resized to have a maximum dimension of 1024 pixels.
-Training is performed on 32 A100 GPUs over five days.
-To ensure training stability, we apply gradient norm clipping with a threshold of 0.1.
-Additionally, we utilize bfloat16 precision and gradient checkpointing to enhance GPU memory and computational efficiency.
-Training Dataset
-For our model training, we utilize approximately 300,000 in-the-wild video sequences collected from public video repositories, along with over 5,173 3D public synthetic static human scans sourced from 2K2K [9], Human4DiT [35], and RenderPeople.
-Specifically, we employ a sampling ratio of 19:1 to draw training batches from the in-the-wild and synthetic datasets to balance generalization and view-consistency.
-To address view bias in the video data, we sample from a diverse range of perspectives as uniformly as possible, guided by the estimated global orientation of SMPL-X.
+2 23.337 0.944 0.091 3.6 min 22.008 0.947 0.067 3.8 min 26.588 0.965 0.036 8.5 min 27.835 0.970 0.018 1.08 sec
+4 23.391 0.943 0.093 6.0 min 23.382 0.952 0.051 5.6 min 27.293 0.966 0.034 15 min 27.940 0.971 0.018 1.40 sec
+8 23.650 0.943 0.090 10.6 min 24.106 0.956 0.048 9.0 min 28.006 0.968 0.031 32 min 28.147 0.971 0.017 2.09 sec
+16 24.235 0.949 0.075 18.8 min 24.398 0.957 0.047 15.7 min 28.358 0.969 0.030 72 min 28.394 0.972 0.016 3.66 sec
+64 25.271 0.956 0.055 54.0 min 25.308 0.960 0.040 66 min 29.587 0.971 0.027 264 min 28.422 0.972 0.016 17.94 sec
+We design three variants of our model with $N_{\text{layer}}=4,6,8$ layers of the PI-MT block, corresponding to LHM++ (S), LHM++ (M), and LHM++ (L), respectively.
+The models comprise approximately 500M, 700M, and 900M training parameters in total.
+The training details can be found in the supplementary material.
+For brevity, LHM++ indicates LHM++ (M) unless otherwise specified.
 Evaluation Protocol
-We report PSNR, SSIM [43], and LPIPS [56] to assess rendering quality, and measure efficiency with GPU memory usage and inference time.
+We report PSNR, SSIM [54], and LPIPS [70] to assess rendering quality.
+If not specifically written, and measure efficiency with GPU memory usage and inference time on a single NVIDIA A100-80G.
+For our evaluation benchmarks, we establish three datasets: public animation videos, smartphone-captured videos for sparse view reconstruction, and in-the-wild fashion videos for both single-view and sparse view settings.
+All evaluation datasets will be made publicly available for future research.
 4.1 Comparison with Existing Methods
 Animatable Human Reconstruction from Sparse Images
-We conduct a comprehensive evaluation of PF-LHM by comparing it with three baseline methods for generating animatable human avatars from casually captured video sequences.
-We assess the efficiency and performance of our model using two types of datasets: one is a public benchmark that includes 20 video sequences from NeuMan [15], REC-MV [30], and Vid2Avatar [7], while the other comprises 24 casual video sequences collected via our smartphones.
-Both Table 2 and Table 3 illustrate quantitative experiments evaluating our model against InstantAvatar [14], GaussianAvatar [11], and ExAvatar [23] on public and casual video sequences.
-Compared to the state-of-the-art (SOTA) baseline ExAvatar, our approach not only significantly accelerates the inference time but also yields comparable quantitative results.
-Specifically, for the model’s efficiency, our model promptly creates animable avatars in seconds while ExAvatar requires approximately 15 minutes to 1.2 hours, depending on the number of input images.
-In terms of model performance, SOTA methods typically require dozens of input images to achieve satisfactory metrics; Moreover, our model achieves more accurate results with substantially fewer inputs, and this capability improves as the number of input images increases.
-As shown in Figure 3, sparse input views lead to noticeable reconstruction artifacts with fitting-based frameworks, including geometric distortions and texture blurring.
-However, our PF-LHM achieves robust and high-fidelity reconstruction from sparse inputs and outperforms the SOTA baselines.
-Table 4: Comparison with single-image method on pose animation using our in-the-wild dataset.
-* indicates we use 80000 query points in LHM for a fair comparison.
-+ denotes the measured inference time of IDOL.
+We first evaluate our method against three state-of-the-art fitting-based approaches designed for sparse inputs: InstantAvatar [17], GaussianAvatar [14], and ExAvatar [29].
+For this evaluation, we utilize two datasets: (1) a public animation dataset comprising 23 video sequences from NeuMan [18], REC-MV [36], and Vid2Avatar [10]; and (2) a smartphone-captured video dataset containing 20 clips collected from in-the-wild environments.
+Figure 3: Animatable human reconstruction comparisons from sparse images.
+Table 3: Comparison with single-image methods on pose animation on our in-the-wild fashion video dataset.
 […]
-AniGS [32] 1 17.023 0.858 0.087 15 minutes 24 GB
-IDOL [60] 1 17.835 0.871 0.083 1.93 seconds + 23 GB
-LHM-0.7B* [31] 1 20.338 0.920 0.047 4.77 seconds 21 GB
-PF-LHM-M 1 20.335 0.919 0.047 1.11 seconds 11 GB
-2 20.891 0.923 0.042 1.42 seconds 11 GB
-4 21.172 0.925 0.039 2.02 seconds 13 GB
-8 21.574 0.926 0.037 3.7 seconds 15 GB
-16 21.752 0.928 0.037 9.15 seconds 18 GB
+AniGS [38] 1 17.234 0.812 0.103 900.00 s 24.00 GB
+IDOL [74] 1 17.912 0.847 0.097 1.93 s 23.12 GB
+LHM-0.7B [37] 1 21.233 0.877 0.078 5.73 s 19.59 GB
+8 21.761 0.883 0.073 24.76 s 25.42 GB
+16 21.868 0.885 0.069 49.55 s 28.77 GB
+64 21.852 0.885 0.069 381.48 s 51.73 GB
+LHM++ 1 21.518 0.882 0.068 0.91 s 6.78 GB
+8 22.208 0.886 0.060 2.09 s 7.35 GB
+16 22.354 0.887 0.057 3.66 s 8.48 GB
+64 22.363 0.887 0.057 17.94 s 29.51 GB
+Table 2 summarizes quantitative experiments evaluating our model against baseline methods on public and smartphone-captured video datasets.
+Specifically, compared to ExAvatar, our method generates animatable avatars within seconds, whereas ExAvatar requires 15 minutes to over an hour.
+Additionally, unlike ExAvatar, which relies on dozens of input images, our model achieves higher accuracy with far fewer inputs.
+As illustrated in Figure 3, while LHM++ marginally outperforms ExAvatar with 16-view inputs, ExAvatar produces severe artifacts such as geometric distortions and texture blurring.
+In contrast, our method yields vivid and realistic animation results.
+Figure 4 shows the avatar reconstruction results of LHM++ in canonical space from sparse images, demonstrating the robustness of the proposed model across diverse input categories, such as real and cartoon-style inputs.
+Figure 4: Human reconstruction of LHM++ in canonical space from sparse image inputs.
+Figure 5: Qualitative comparison with LHM.
+LHM++ matches LHM’s quality with a single input view and generates progressively more detailed results as view count increases.
+Please zoom in for better view.
 Animatable Human Reconstruction from a Single Image
-We evaluate PF-LHM against three baseline approaches for single-view animatable human reconstruction.
-The first baseline is AniGS [32], which employs a multi-view diffusion model to create canonical human avatars, followed by 4D Gaussian splatting (4DGS) optimization to address inconsistencies across different views.
-The second one is IDOL [60], using a UV-based transformer model to create the animatable avatars training from synthetic human datasets built on a human video diffusion model.
-The last baseline, LHM [31], introduces novel body-head transformer blocks that directly regress human Gaussian parameters in canonical space.
-For our evaluation, we employ 400 in-the-wild video sequences featuring individuals of various age groups, including young men and women, older adults, and children.
-For a fair comparison, we compare LHM with the same parameters and query points.
-Figure 7 shows a qualitative comparison experiments with LHM on in-the-wild data.
-The figure indicates that PF-LHM achieves results comparable to LHM with single-view image input.
-Furthermore, as the number of inputs increases, our method generates increasingly realistic and detailed results.
-Table 4 presents two key findings regarding single-view human reconstruction.
-Our unified framework not only achieves competitive quantitative results compared to LHM but also the inference speed that is four times faster than that of LHM.
-Furthermore, as the number of input views increases, our method’s performance also improves.
-Specifically, in comparison to a single-image input, using 16 input views results in improvement of 1.417, 0.09, and 0.01 in PSNR, SSIM, and LPIPS, respectively.
-Figure 7 shows a qualitative comparison experiment with LHM.
-4.2 Qualitative Results
+To assess performance under extreme sparse conditions, single-view input, we compare LHM++ against two state-of-the-art single-image methods: AniGS [38] and IDOL [74].
+We use 400 in-the-wild fashion videos as the evaluation benchmark.
+As demonstrated in Table 3, LHM++ outperforms AniGS and IDOL by $4.28$ and $3.61$ dB in PSNR, respectively.
+In terms of LPIPS, we observe improvements of $0.35$ and $0.29$ against the same baselines, validating the superiority of our approach in single-view settings.
+Animatable Human Reconstruction from Any Images
+We also compare against LHM [37], which was originally trained on single-view data.
+To enable it to handle arbitrary-view inputs, we align its configuration with our method and fine-tune it for any-view inputs.
+Notably, we retain LHM’s original configuration of 40K query points, as increasing this number to match our method resulted in out-of-memory errors.
+As shown in Table 3, LHM++ consistently outperforms LHM across 1-, 8-, 16-, and 64-view inputs, achieving PSNR gains of $0.285$, $0.447$, $0.486$, and $0.511$ dB, respectively.
+Beyond accuracy, our method achieves approximately $10\times$ faster inference speed while consuming only a quarter of the memory compared to LHM.
+Figure 5 presents a qualitative comparison using 1- and 8-view inputs, where our method produces significantly more realistic and detailed results, demonstrating robust scalability from single to dense views.
+Additional qualitative and quantitative results can be found in the appendix.
+Table 4: Effectiveness of the Encoder-Decoder PIT under varying inputs and points.
+# Points 1 view 4 views 8 views 16 views
+LHM-0.7B LHM++ LHM-0.7B LHM++ LHM-0.7B LHM++ LHM-0.7B LHM++
+40 K 5.16 s 0.55 s 9.41 s 0.74 s 15.23 s 1.01 s 32.32 s 1.74 s
+80 K 19.47 s 0.69 s 25.73 s 0.88 s 35.36 s 1.19 s 59.25 s 1.95 s
+160 K 74.31 s 0.79 s 85.58 s 1.00 s 102.44 s 1.31 s 140.67 s 2.13 s
+4.2 Ablation Study
+In this section, we present comprehensive ablation studies demonstrating that LHM++ is not a simple extension of LHM but a carefully designed, novel, and efficient encoder–decoder framework.
+Figure 6: The ablation study for both (a) tokens compression and (b) 3D-aware human neural renderer.
+‘NR’ is the abbreviation for ‘3D-aware neural renderer’.
+The red boxes highlight the visual difference.
+Please zoom in for better view.
+Effectiveness of the Encoder-Decoder PIT
+[…]
+Token Compression Across Image Numbers
+Our framework accelerates attention process by compressing 3D geometric tokens via grid pooling and 2D image tokens through token merging (details can be found in supplementary).
+To evaluate the efficacy of these strategies, Fig.
+6(a) analyzes the impact of token compression across varying numbers of input views.
+Under sparse-view settings, geometric tokens dominate inference cost; compressing them yields major speedups.
+As the number of input image increases, image tokens become predominant, shifting the computational bottleneck to image token processing.
+By jointly compressing both token types, our method ensures consistent efficiency gains across both sparse and dense input configurations.
+Table 5: Comparison with LHM for multi-modality fusion on our in-the-wild fashion dataset with 8 input views.
+Methods 40 K 80 K 160 K
+PSNR Time PSNR Time PSNR Time
+LHM-0.7B 21.761 24.76 s 21.803 49.78 s 21.796 110.37 s
+LHM++ 21.747 1.86 s 22.124 1.98 s 22.208 2.09 s
+Effect of 3D Query Token Count
+Building on the analysis presented above, we next evaluate the performance improvements of the proposed method over LHM.
+First, we study how the number of 3D query tokens influences both methods.
+Table 5 reports comparative experiments on the in-the-wild fashion dataset with eight input views, varying the number of sampling points.
+The results indicate that LHM gains only marginally as the count of 3D tokens increases.
+We attribute this to LHM’s inefficient fusion mechanism.
+LHM’s architecture learns very large attention maps, so redundant tokens impede learning and degrade fusion efficiency.
+By contrast, the encoder–decoder PIT design effectively prunes redundant 3D and 2D tokens, substantially improving the model’s fusion performance.
+Effect of Attention Modules
+We conduct an ablation study to investigate the impact of attention modules within the encoder–decoder PIT architecture on multi-modal fusion across in-the-wild fashion, as well as public animation and smartphone-captured video datasets (Public & Smartphone).
+As detailed in Table 6, incorporating image-wise attention enhances per-frame feature representation compared to relying solely on multi-modal attention.
+Furthermore, point-wise attention facilitates the fusion of local geometric features by leveraging Euclidean distances within the point cloud.
+This mechanism promotes effective point-to-point context exchange, also improving overall model performance.
+Finally, the synergistic combination of both attention modules yields the most significant performance gains.
+Table 6: effect of attention modules in the encoder–decoder PIT.
+Methods In-the-wild Fashion Public & Smartphone
+4 Views 16 Views 4 Views 16 Views
+LHM++ w/o Image/Point-wise Attention 21.669 21.977 27.403 27.655
+LHM++ w/o Image-wise Attention 21.735 22.136 27.656 27.927
+LHM++ w/o Point-wise Attention 21.838 22.230 27.764 28.129
+LHM++ 21.957 22.354 27.940 28.394
+Effect of 3D-Aware Human Neural Renderer
+Figure 6(b) illustrates the qualitative results of the ablation study for 3D-aware neural renderer.
+Since our LHM++ directly learns a positional residual from the SMPL-X template, challenges remain in rendering loose-fitting clothing, despite employing mask distribution loss.
+The proposed neural renderer not only effectively addresses this issue, but also significantly enhance the overall rendering results.
+Additional ablation studies are provided in the appendix, including scalability, effect of image token merging and the detailed analysis on neural renderer.
+5 Conclusion
+We present LHM++, an efficient feed-forward framework for rapid and high-fidelity 3D human avatar reconstruction from one or a few casually captured, pose-free images.
+Our approach introduces the Encoder-Decoder Point-Image Transformer (PIT), which enables efficient multimodal fusion between geometric point tokens and sparse multi-view image tokens through hierarchical attention.
+In addition, we introduce a lightweight 3D-aware neural animation renderer to refine the rendering quality of reconstructed avatars.
+Extensive experiments across synthetic and real-world datasets demonstrate that LHM++ effectively unifies single- and sparse-view reconstruction and supports realistic avatar animation.
+Appendix A Overview
+This supplementary material provides comprehensive details supporting the main paper.
+We begin with Sec. B (Demo Video), which presents animation results, comparative experiments, and a visual gallery across diverse datasets.
+Subsequently, Sec.
+C (Method Details and Analysis) encompasses the Encoder-Decoder Point-Image Transformer architecture (Sec.
+C.1), image token merging strategies (Sec.
+C.2), and Gaussian regularization losses (Sec.
+C.3).
+This section also provides a detailed analysis of the 3D-aware neural renderer (Sec.
+C.4) and our implementation configurations (Sec.
+C.5).
+Next, Sec.
+D (Experiments) provides comprehensive experimental details, including the data protocols employed from public animation datasets (Sec.
+D.1).
+It presents comparisons with commercial image-to-3D methods (e.g., Hunyuan-2.5) (Sec.
+D.2) and extensive ablation studies on scalability (Sec.
+D.3) and training/inference efficiency (Sec.
+D.4).
+Furthermore, we provide comprehensive quantitative (Sec.
+D.5) and qualitative results (Sec.
+D.6) on evaluation datasets to demonstrate the model’s robustness and generalization capabilities.
+Finally, Sec. E discusses the limitations of LHM++, presents representative failure cases, and proposes potential remedies.
+Appendix B Demo Video
+Please refer to the Demo Video for animation results of the reconstructed 3D avatars.
+Our demo consists of three components: (1) rendering results from varying input views, (2) comparison experiments against methods such as ExAvatar [29], LHM [37], and IDOL [74], and (3) an animated video gallery across public animation and smartphone-captured datasets, in-the-wild fashion videos, and AI-generated images.
+Details of Testing Dataset in Demo
+For comparison experiments with ExAvatar, we utilize 16-view images sampled uniformly from 00000_random and 00069_Dance in Vid2Avatar [11], anran_self_rotated from REC-MV [36], and Bike from NeuMan [18].
+For comparison experiments with feed-forward approaches, we use Anran_Skirt from REC-MV [36] and two in-the-wild monocular video samples.
+Next, we provide animated gallery videos to evaluate our model’s generalization capability across four public datasets.
+For REC-MV [36], we select anran_purple and xiaolin.
+For NeuMan [18], we use clips from bike, citron, and jogging.
+For Vid2Avatar [11], we choose Yuliang, 00000_random, exstrimalik, and Suarez.
+Lastly, for MVHumanNet [60], we utilize samples with IDs 101157, 103383, 103528, and 104316.
+Details of In-the-Wild Fashion Videos
+For in-the-wild fashion videos, we utilize web-sourced monocular videos and randomly sample 8 views from the original clips.
+Details of AI-Generated Images
+For AI-generated images, we employ FLUX.1 Kontext [4] to generate four different viewpoint images using prompts such as “the front view of”, “the back view of”, “the right view of”, and “the left view of”.
+Appendix C More Details of Method and Analysis
+C.1 Encoder-Decoder Point-Image Transformer
+Points Serialization
+To ensure both scalability and efficiency within our feed-forward framework, we employ serialization to convert unstructured SMPL-X anchor points into a structured format.
+Following Point Transformer v3 (PTv3) [59], we utilize a mixture of four serialization patterns: Z-order, Hilbert, Trans Z-order, and Trans Hilbert.
+To prevent the model from overfitting to a specific point ordering, we randomly shuffle the serialization pattern sequence order to each PTv3 block.
+Figure 7: Detailed architecture of the PTv3 Block [59].
+PTv3 Block
+As depicted in Fig.
+7, our point attention module adopts the architecture from Point Transformer v3 [59], leveraging patch-based self-attention to accelerate inference.
+Additionally, we incorporate Grid Pooling to downsample the point cloud, which significantly improves the computational efficiency of the self-attention mechanism.
+Regarding implementation details, patch sizes are configured to decrease progressively across downsampling stages.
+Specifically, we use patch sizes of {4096, 2048, 1024} for 80K point clouds and {8192, 4096, 2048} for 160K point clouds.
+Figure 8: Illustration of the image token merging process.
+Given 2D tokens $\mathbf{T}_{\text{2D}}$ and key features $\mathbf{K}_{\text{2D}}$ from the preceding frame-wise attention block, source tokens are randomly sampled.
+A similarity matrix is then computed between the source and target tokens (derived from $\mathbf{K}_{\text{2D}}$), allowing the most similar source tokens to be greedily merged into their corresponding target tokens.
+3D Token Evolution
+Table 7 reports the evolution of 3D token counts across the six encoder layers of the PIT.
+To manage this reduction, we apply a point-cloud grid-pooling operator that efficiently compresses redundant tokens based on their Euclidean proximity.
+Table 7: Number of 3D tokens across the encoder layers of the PIT.
+Operator 40K 80K 160K
+Grid Pooling 1 16778 22380 24377
+Point-Image Attention 1 16778 22380 24377
+Grid Pooling 2 5547 6036 6441
+Point-Image Attention 2 5547 6036 6441
+Grid Pooling 3 1493 1625 1686
+Point-Image Attention 3 1493 1625 1686
+C.2 Image Token Merging
+As the number of input images increases, the image tokens gradually dominate the computing resources.
+Inspired by ToME [5, 6], we employ Bipartite Soft Matching to merge redundant image tokens, which helps to reduce inference times during the global attention process.
+As illustrated in Fig.
+8, we construct a bipartite graph using stride sampling as proposed in ToME-SD [6].
+To build the similarity matrix, we utilize the ’Key’ tokens from the frame-wise attention mechanism described in the paper.
+Subsequently, we merge the similar image tokens based on a fixed merge ratio $r$.
+The merged image tokens are then inputted to multi-modal attention with the geometric 3D tokens.
+After the multi-modal attention operation, we unmerge the image tokens back to their original size using a simple feature copying operator, allowing them to proceed to the next PI transformer block, and not change their original size.
+The effect of Image Token Compression
+Surprisingly, we found that reducing a certain proportion of image tokens does not degrade model performance.
+in contrast, it can even lead to a slight improvement.
+We hypothesize that the performance gain benefits from the removal of redundant tokens, making the attention map learning more effective.
+Table 8 presents the details of the ablation study examining the ratio of token merging.
+Table 8: Effect of image token merging across reduction ratios The metrics are calculated based on 16-view image inputs on in-the-wild fashion videos dataset.
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$
+ratio=0.00 22.317 0.886 0.059 5.73 s
+ratio=0.25 22.352 0.887 0.058 4.82 s
+ratio=0.50 22.354 0.887 0.057 3.66 s
+ratio=0.75 22.335 0.886 0.059 3.07 s
+Analysis on Image Token Merging
+To further analyze the ‘merge’ and ‘unmerge’ operator in global image-point attention, we conducted both qualitative and quantitative experiments that demonstrate how performance varies with different merge ratios and numbers of views.
+Figure 9 illustrates qualitative results showing how visual quality changes with different token compression ratios across sparse-view inputs.
+The figure indicates that for $1\sim 4$ views a high merge ratio $r>0.5$ significantly degrades visual quality, whereas for $8\sim 16$ views, larger merge ratios $r\in[0.5,0.75]$ do not obviously reduce performance.
+We attribute this to the increased redundancy of tokens as the number of views grows, and the our model de facto does not require all tokens for effective reconstruction.
+Moreover, Table 9 presents a quantitative ablation study of image-token merging.
+The table reports PSNR averaged across both public animation and smartphone-captured video datasets.
+the result shows that applying token merging with moderate ratios $r\in[0.25,0.75]$ for sparse-view inputs $4\sim 16$ yields metrics comparable to, or slightly better than, the no-merge baseline $r=0$.
+These results are consistent with findings from previous work in image generation [6].
+Figure 9: Ablation study of image-token merging.
+‘r’ denotes the merge ratio ($r=0$ means no merging).
+Larger ‘r’ values correspond to greater compression of image tokens.
+Please zoom in for better view.
+Table 9: Ablation study of image-token merging. The PSNR metrics are averages computed over public and smartphone-captured video sequences.
+Input Views $r=0.00$ $r=0.25$ $r=0.50$ $r=0.75$ $r=0.90$ $r=0.95$
+view = 1 27.762 27.735 27.733 27.644 27.525 27.238
+view = 4 27.942 27.942 27.940 27.937 27.893 27.662
+view = 8 28.140 28.153 28.147 28.135 27.903 27.735
+view = 16 28.391 28.394 28.394 28.391 28.358 28.206
+C.3 Gaussian Regularization Loss
+Geometric Regularization
+Video-based supervision frequently suffers from incomplete coverage of the human body and typically exhibits limited motion diversity.
+These limitations often lead to geometric artifacts and ambiguities in unobserved regions.
+To mitigate these issues, we incorporate two geometric regularizers adopted from LHM [37]: (1) the As Spherical As Possible loss, $\mathcal{L}_{\text{ASAP}}$, which enforces isotropy in the 3D Gaussians; and (2) the As Close As Possible loss, $\mathcal{L}_{\text{ACAP}}$, which preserves spatial coherence among neighboring primitives.
+Mask Distribution Regularization
+Furthermore, empirical observations indicate that the optimization process tends to favor expanding Gaussian scales over refining positional offsets to represent geometry.
+This behavior induces significant artifacts, particularly when animating avatars with loose clothing in dynamic poses.
+To counteract this tendency, we propose a Mask Distribution Loss, $\mathcal{L}_{\text{dis}}$, which encourages Gaussian primitives to prioritize positional accuracy over scale expansion.
+Specifically, this regularization is implemented by rendering an auxiliary mask, $M_{\text{dis}}$, using fixed Gaussian attributes (opacity $\rho=0.95$, scale $\sigma=0.002$), and minimizing the L1 discrepancy between $M_{\text{dis}}$ and the ground-truth human mask.
+By fixing the scale and opacity, this loss underscores the contribution of Gaussian positions.
 As demonstrated in Fig.
-8, we present the animation results of avatars generated from various in-the-wild monocular videos, including NeuMan [15], REC-MV [30], Vid2Avatar [7], MVHumanNet [49] and our causal video dataset.
-Our PF-LHM is capable of generalizing across different identities and garment styles, producing highly realistic renderings for novel human poses and arbitrary viewpoints.
-Figure 4: Animatable human reconstruction comparisons from sparse images on in-the-wild videos.
-PF-LHM-S PF-LHM-M PF-LHM-L GT
-Figure 5: Ablation study on model design and parameters.
-Figure 6: PSNR performance of our models with different input views.
-Figure 7: Comparison results of animatable human reconstruction methods with LHM.
-Figure 8: Visual animation results of avatars created from monocular in-the-wild videos.
-The created 3D avatars can be animated using novel human poses and demonstrate highly detailed appearance.
-Also, the figure effectively illustrates the generalization capability of our approach, delivering outstanding results across various demographics, including the elderly, children, men, and women, as well as individuals of differing heights and body types.
-Table 5: Model Efficiency.
-We evaluate the training and inference efficiency of various backbones.
-The batch size is set to 1, the number of input views is equal to 16.
-‘# Points’ refers to the number of geometric points, and ‘Time’ denotes the duration of a single iteration during both training and inference.
+10, incorporating this term encourages a more uniform distribution of Gaussian centers across the ground-truth mask area, effectively preventing the model from converging to solutions with excessively large scale parameters.
+Input w/o $\mathcal{L}_{\text{dis}}$ w/ $\mathcal{L}_{\text{dis}}$ Human Mask
+Figure 10: Visual ablation study of the mask distribution loss.
+The comparison illustrates how $\mathcal{L}_{\text{dis}}$ promotes a more uniform spatial distribution of Gaussian primitives, aligning better with the ground-truth human mask.
+The combined geometric regularization term is formulated as:
+$\mathcal{L}_{\text{reg}}=\lambda_{\text{dis}}\mathcal{L}_{\text{dis}}+\lambda_{\text{ASAP}}\mathcal{L}_{\text{ASAP}}+\lambda_{\text{ACAP}}\mathcal{L}_{\text{ACAP}},$ (12)
+where the weighting coefficients are empirically set to $\lambda_{\text{dis}}=0.5$, $\lambda_{\text{ASAP}}=20$, and $\lambda_{\text{ACAP}}=5$.
+C.4 3D-Aware Human Animation Neural Renderer
+To improve the rendering quality, we develop a lightweight neural renderer built on the 3DGS renderer to refine the final output.
+Specifically, we utilize the human prior model SMPL-X to animate the reconstructed 3D avatar to the target pose, then render high-dimension semantic features via 3DGS rasterization.
+Subsequently, we predict final render results using the lightweight neural renderer.
+Rendering Time Across Different Resolutions
+The entire rendering process can be run in real time. Table 10 clearly illustrates the rendering time of this process across different resolutions.
+Table 10: Rendering Time of the 3D-aware neural render across different resolutions on a Single NVIDIA A100-80G
+Resolution Time (s) $\downarrow$ FPS $\uparrow$
+$592\times 592$ 0.013 78
+$880\times 880$ 0.027 37
+$1176\times 1176$ 0.048 21
+Effect of 3DGS Representation for Neural Renderer
+Our renderer consists of a Gaussian Splatting rasterizer and a lightweight neural renderer.
+This raises an important question: Is the 3DGS intermediate representation necessary?
+Could the framework rely solely on a neural renderer, similar to HumanRAM [67]?
+To answer this question, we performed a comparison experiment to validate our hypothesis: the 3DGS representation plays a crucial role in the neural renderer.
+For the without-3DGS experiments, we employ point rasterization instead of the 3DGS rasterizer.
+Specifically, after the encoder-decoder PIT, we deform the points with high-dimensional semantic features and render point features via point rasterization implemented with PyTorch3D.
+Then, we use the neural renderer to obtain the final results.
+For training, apart from the 3DGS representation, all settings were identical to LHM++.
+As shown in Fig.
+11, the comparison experiment illustrates that relying solely on point offsets without an explicit 3DGS representation causes the neural renderer to struggle with reproducing the input-image context, producing high-frequency artifacts and blurring.
+Moreover, the pipeline experiences convergence difficulties.
+In contrast, employing an explicit 3DGS representation enables faster convergence of the neural renderer and substantially improves overall visual quality.
+w/o GS w/ GS GT w/o GS w/ GS GT w/o GS w/ GS GT
+Figure 11: The effects of 3DGS representation for neural renderer. ‘GS’ indicates 3D Gaussian Splatting.
+C.5 Implementation Details
+Training Strategy
+To ensure training stability, we initially train LHM++ for single-view input by minimizing the training loss using the AdamW optimizer for 65 K iterations, excluding the 3D-aware neural renderer.
+Subsequently, to enable a pose-free setting capable of handling arbitrary image inputs, we randomly sample 1 to 16 frames from a training video for each batch.
+We then continue training for an additional 41 K iterations using the same optimizer and loss objectives.
+We apply gradient norm clipping with a threshold of 0.1 and employ bfloat16 precision with gradient checkpointing to optimize GPU memory usage and computational efficiency.
+A cosine learning rate scheduler is employed with a peak learning rate of $1\text{e-}4$ and a warm-up period of 3,000 iterations.
+Input images are resized to a maximum dimension of 1024 pixels.
+Training is conducted on 32 A100 GPUs for approximately one week.
+Training Dataset
+For training, we utilize approximately 300,000 in-the-wild video sequences collected from public repositories, complemented by over 5,173 synthetic static 3D human scans sourced from 2K2K [12], Human4DiT [42], and RenderPeople.
+Specifically, we sample training batches from the in-the-wild and synthetic datasets at a ratio of 19:1 to balance generalization and view consistency.
+To mitigate view bias in the video data, we sample perspectives uniformly across a diverse range, guided by the estimated global orientation of SMPL-X.
+Appendix D Experiments
+D.1 Details of Public Animation Video
+We collect 23 public videos to form our evaluation benchmark.
+Specifically, we selected anran_dance_self_rotated, anran_purple, anran_skirt, lingteng, self-rotate-leyang, and xiaolin from Rec-MV [36].
+Additionally, we sampled bike, citron, parkinglot, jogging, lab, and seattle from NeuMan [18].
+Furthermore, we included 00000_random, 00020_Dance, 00069_Dance, exstrimalik, helge, lam, Nadia, phonecall, roger, truman, and Yuliang from Vid2Avatar [10, 11].
+Recon Reset Recon Reset Recon Reset
+Figure 12: Comparison with Hunyuan3D-2.5.
+We use Hunyuan3D-2.5 to reconstruct 3D models in the input-view space.
+We then auto-rig the models using Mixamo and reset the target-view pose into the canonical space.
+”Recon” denotes reconstruction in the input space, while ”Reset” denotes animating the avatar into a predefined canonical pose.
+D.2 Comparison with General Image-to-3D Generation
+We conducted comparative experiments against state-of-the-art general image-to-3D generation methods.
+These include cutting-edge commercial software, such as Rodin [69] and Hunyuan3D-2.5 [48].
+For our experiments, we selected Hunyuan3D-2.5 as the baseline due to its state-of-the-art performance.
+Since Hunyuan3D only reconstructs static 3D assets, we use Mixamo to auto-rig the reconstructed models after 3D avatar reconstruction, making reconstructed static 3D avatar animatable.
+As shown in Fig.
+12, successful rigging of a 3D digital avatar demands stringent input pose requirements.
+Although Hunyuan3D-2.5 can recover the overall structure from the input image, it often fails to produce an ideal riggable geometry.
+This is evident in the ”Reset” panel, where anatomical points on the arms and torso are difficult to distinguish, complicating avatar animation.
+Furthermore, the generated textures often appear unnatural, containing artifacts and unrealistic details.
+In contrast, LHM++ generates natural, high-fidelity avatars in canonical space with superior rigging compatibility.
+D.3 Scalability Analysis
+Dataset Scalability
+To assess dataset scalability, we performed controlled experiments using stratified random subsets of 10K and 100K videos from the original 300K training dataset.
+Table 11 demonstrates that relying solely on synthetic data leads to poor model generalization.
+In contrast, incorporating in-the-wild data significantly improves model robustness and performance on real-world evaluations.
+Furthermore, increasing the dataset size yields consistent performance gains.
+Figure 13 illustrates the visual quality improvements as the number of in-the-wild videos increases.
+Synthetic 10K 100K 300K GT
+Figure 13: Ablation study on dataset scalability.
+Model Parameter Scalability
+To verify the scalability of our LHM++, we train variant models with increasing parameter numbers by scaling the layer numbers.
+Our experiments indicate that increasing the number of model parameters correlates with improved performance.
+However, the performance improvements tend to saturate as the model size continues to grow.
+We believe that this limitation is largely influenced by the scale of our training dataset that is extremely smaller—by an order of magnitude—than the extensive datasets employed in LLMs.
+Our experiments reveal that while performance shows moderate enhancements from small to medium-sized models, the gains become negligible for larger models.
+Table 11: Ablation study on dataset scalability. The metrics are calculated based on 16-view image inputs on the in-the-wild video dataset.
+Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+Synthetic Data 20.549 0.858 0.086
+10K Videos 21.526 0.870 0.073
+100K Videos 22.051 0.885 0.062
+300K Videos + Synthetic Data 22.354 0.887 0.057
+D.4 Training and Inference Efficiency
+[…]
+Table 12: End-to-end efficiency. Batch size=2, input views=8. # Points = geometric points; Time = per-iteration duration (training/inference).
 Methods Params. # Points Training Inference
 Time Memory Time Memory
-LHM-0.5 B [31] 500 MB 80 K 26.33 s 54.97 GB 21.61 s 27.8 GB
-LHM-0.7 B [31] 700 MB 80 K 50.23 s 62.27 GB 43.06 s 28.3 GB
-LHM-1.0 B [31] 1000 MB 80 K 70.61 s 71.55 GB 64.29 s 29.7 GB
-LHM-1.0 B [31] 1000 MB 160 K - $>$ 80 GB 159.19 s 35.8 GB
-PF-LHM-S 500 MB 80 K 8.77 s 46.5 GB 7.35 s 17.0 GB
-PF-LHM-M 700 MB 80 K 10.01 s 50.4 GB 9.03 s 18.3 GB
-PF-LHM-L 1000 MB 80 K 11.13 s 55.1 GB 10.11 s 19.2 GB
-PF-LHM-L 1000 MB 160 K 11.20 s 55.8 GB 10.64 s 20.0 GB
-4.3 Ablation Study
-Model Efficiency
-Table 5 presents quantitative results that substantiate the efficiency of our model, tested on NVIDIA A100-80G hardware.
-The table clearly illustrates that our novel framework significantly outperforms the original LHM architecture, achieving notably reduced training times and lower memory consumption.
-Furthermore, during testing, our inference times are about 5$\sim$10 times faster than those of LHM.
-Model Parameter Scalability
-To verify the scalability of our PF-LHM, we train variant models with increasing parameter numbers by scaling the layer numbers.
-Table 6 compares performance across various model capacities.
-Our experiments indicate that increasing the number of model parameters correlates with improved performance.
-Figure 6 presents the comparison among PF-LHM-S, PF-LHM-M, and PF-LHM-L where the larger model achieves more accurate reconstruction.
-Also, Fig.
-6 depicts the performance of different models based on varying input view counts, clearly indicating that our model is scalable and that performance improves with an increased number of input images.
-Number of Query Points
-Table 6 shows an ablation study analyzing the effect of varying the number of query points on public video datasets.
-As the number of query points increases from 40K to 80K, our model demonstrates improvements in PSNR, SSIM, and LPIPS by 0.562, 0.006, and 0.002, respectively.
-However, when the number of query points is increased further from 80K to 160K, we observe a slight gain in performance.
-Therefore, we set the number of query points to 80K to achieve an optimal balance between efficiency and model performance.
-Table 6: Analysis of model parameters and 3D geometric point numbers.
-[…]
-PF-LHM-M 40 K 28.001 0.969 0.017 7.32 seconds 17 GB
-PF-LHM-M 160 K 28.565 0.975 0.014 10.64 seconds 20 GB
-PF-LHM-S 80 K 28.089 0.970 0.017 7.35 seconds 17 GB
-PF-LHM-M 80 K 28.563 0.975 0.015 9.03 seconds 18 GB
-PF-LHM-L 80 K 28.924 0.977 0.014 10.11 seconds 19 GB
-5 Conclusion
-We present PF-LHM, a novel feed-forward framework for rapid and high-fidelity 3D human avatar reconstruction from one or a few casually captured, pose-free images.
-Our approach introduces the Encoder-Decoder Point-Image Transformer (PIT), which enables efficient multimodal fusion between geometric point tokens and sparse multi-view image tokens through hierarchical attention.
-By leveraging the PIT architecture, our method achieves superior scalability, generalization, and reconstruction quality while significantly reducing runtime compared to optimization-based baselines.
-Extensive experiments across synthetic and real-world datasets demonstrate that PF-LHM effectively unifies single- and sparse-view reconstruction and supports realistic avatar animation using 3D Gaussian Splatting.
-Appendix A Implementation Details
-A.1 Details of Points Serialization
-To trade the scalability and efficiency of our feed-forward framework, we leverage serialization to transform unstructured SMPL-X anchor points into structured data format.
-Following Point Transformer v3 (PTv3) [48], we mixture 4 patterns of serialization, including Z-order, Hibert, Trans Z-order and Trans Hibert, and apply random shuffle to the order of serialization patterns before each PTv3 block.
-Figure 9: Detailed architecture of PTv3 Block [48].
-A.2 Details of PTv3 Block
-As illustrated in Fig.
-9, our point attention architecture integrates the mechanism introduced in Point Transformer v3 [48].
-This architecture employs patch-based self-attention to accelerate the forward pass.
-Furthermore, Grip Pooling is utilized to downsample the point cloud, enhancing the efficiency of point cloud self-attention.
-Importantly, to facilitate interaction among different patch groups, a shuffling order is implemented.
-In our setup, the patch sizes are set to 4096, 2048, and 1024 for the 80K point clouds, decreasing in accordance with the downsampling process.
-For the 160K point clouds, the patch sizes are configured to 8192, 4096, and 2048.
-Table 7: Ablation study on dataset scalability. The metrics are calculated based on 16-view image inputs.
-Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-PF-LHM-L + Synthetic Data 26.148 0.952 0.027
-PF-LHM-L + 10K Videos 27.456 0.970 0.018
-PF-LHM-L + 100K Videos 28.312 0.975 0.017
-PF-LHM-L + All 28.924 0.977 0.014
-Only Synthetic 10K Videos 100K Videos 300K Videos GT
-Figure 10: Ablation study on dataset scalability.
-Ground Truth w/o Deformation w/ Deformation
-Figure 11: The ablation study for posed-conditioned deformation. The red boxes highlight the visual difference mentioned above.
-Input w/o $\mathcal{L}_{\text{dis}}$ w/ $\mathcal{L}_{\text{dis}}$ Human Mask
-Figure 12: The ablation study for mask distribution loss.
-Appendix B Experiments
-B.1 Details of Testing Dataset
-For our qualitative results, we utilize videos from Vid2Avatar [7], specifically those indexed as: 00000_random, 00020_Dance, 00069_Dance, exstrimalik, and Yuliang.
-Additionally, we draw on videos from Rec-MV [30], including the following indices: anran_dance_self_rotated, anran_purple, anran_skirt, self-rotate-leyang, and xiaolin.
-Moreover, we use the NeuMan [15] dataset with the indices: citron and seattle.
-And, sparse view images are randomly sampled from MVHumanNet [49], with the following indices: 101330, 101425, 101483, 103383, 103528, 104013, 203243, and 103594, along with videos captured using our smartphone.
-B.2 More Ablation Study
-Details of Dataset Scalability
-To assess the scalability of our dataset, we perform controlled experiments utilizing stratified random subsets of 10K and 100K from the original training dataset of 300K videos.
-Table 7 demonstrates that relying solely on the synthetic dataset leads to poor model generalization.
-In contrast, incorporating an in-the-wild dataset significantly improves the model’s robustness and performance during real-world evaluations.
-Furthermore, increasing the size of the dataset siginificantly improves model results, although the gains in performance tend to diminish with larger dataset sizes.
-Figure 10 illustrates the findings of our ablation study regarding dataset scalability.
-Mask Distribution Loss
-Gaussian models tend to learn large-scale Gaussian primitives rather than large step offsets.
-To address this issue, we introduce a mask distribution regularization loss that encourages Gaussian primitives to learn offsets instead of concentrating on large-scale axes.
-As illustrated in Fig.
-12, employing this loss enables the mean position of the Gaussian primitives to be distributed more evenly across the ground truth mask area, thereby preventing the model from learning large-scale axes of the Gaussian primitives.
-Posed-Condition Deformation
-Figure 11 illustrates the impact of using pose-conditioned deformation.
-The results clearly indicate that our deformation tokens, which are conditioned on SMPL-X parameters, effectively model non-rigid deformations of loose garments, in contrast to the case without pose conditioning.
-B.3 More Results
-Figure 13 presents more animation results using eight input images.
+LHM-0.7B* [37] 700 M 40 K 81.22 s 80.3 GB 24.76 s 25.42 GB
+700 M 160 K - OOM 110.3 s 36.71 GB
+LHM++ Small 500 M 160 K 5.79 s 43.5 GB 1.58 s 6.98 GB
+LHM++ Medium 700 M 160 K 6.35 s 48.7 GB 2.09 s 7.35 GB
+LHM++ Large 900 M 160 K 7.25 s 52.6 GB 2.88 s 8.21 GB
+D.5 More Quantitative Results
+Table 3 of the main paper summarizes the averaged results on the public and smartphone-captured video sequences. Table 13 and Table 14 show the detailed results for each dataset.
+Table 13: Comparison experiments with sparse-view input methods on public animation dataset.
+views InstantAvatar [17] GaussianAvatar [14] ExAvatar [29] LHM++ M
+PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time
+2 21.718 0.929 0.125 3.6 m 21.049 0.937 0.086 3.8 m 26.371 0.959 0.045 8.5 m 26.900 0.964 0.022 1.08 s
+4 21.782 0.926 0.126 6.0 m 22.614 0.943 0.065 5.6 m 26.906 0.960 0.044 15 m 27.031 0.965 0.022 1.40 s
+8 22.308 0.928 0.120 10.6 m 23.238 0.947 0.061 9.0 m 27.319 0.962 0.040 32 m 27.346 0.965 0.021 2.09 s
+16 23.026 0.936 0.101 18.8 m 23.427 0.949 0.061 15.7 m 27.456 0.962 0.040 1.2 h 27.521 0.966 0.020 3.66 s
+64 23.561 0.944 0.077 54.0 m 23.574 0.949 0.055 1.1 h 27.896 0.962 0.037 4.4 h 27.557 0.966 0.020 17.94 s
+Table 14: Comparison experiments with sparse-view input methods on smartphone-captured videos.
+views InstantAvatar [17] GaussianAvatar [14] ExAvatar [29] LHM++ M
+PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time PSNR SSIM LPIPS Time
+2 24.340 0.954 0.070 3.6 m 22.602 0.953 0.056 3.8 m 26.723 0.969 0.031 8.5 m 28.413 0.974 0.016 1.08 s
+4 24.387 0.953 0.072 6.0 m 23.857 0.957 0.043 5.6 m 27.532 0.970 0.028 15 m 28.502 0.974 0.016 1.40 s
+8 24.481 0.952 0.072 10.6 m 24.644 0.961 0.040 9.0 m 28.431 0.972 0.026 32 m 28.643 0.974 0.015 2.09 s
+16 24.984 0.957 0.059 18.8 m 24.999 0.962 0.039 15.7 m 28.916 0.973 0.024 1.2 h 28.934 0.975 0.014 3.66 s
+64 26.329 0.964 0.042 54.0 m 26.381 0.967 0.030 1.1 h 30.634 0.978 0.020 4.4 h 28.958 0.975 0.013 17.94 s
+D.6 More Qualitative Results
+Human Reconstruction in Canonical Space
+Figure 15 demonstrates human reconstructions in canonical space from sparse views, showing that our model can reconstruct reasonable, high-quality avatars from sparse images input.
+More Animation Rendering Results
+Both Fig.
+16 and Fig.
+17 present more animation results using eight input images.
 Our method enables high-fidelity reconstruction and animation of human avatars through the efficient Point-Image Transformer architecture, thereby demonstrating robust generalization capabilities and practical effectiveness.
-Appendix C Limitations and Future Work
-A primary limitation of our approach lies in its reliance on the SMPL-X template mesh for initializing geometric tokens.
-While this offers a strong structural prior, it can constrain reconstruction fidelity for subjects wearing loose or non-body-conforming garments, such as dresses, which deviate significantly from the SMPL-X topology.
-Additionally, due to the limited diversity of large motion poses in our training datasets, the model’s performance may degrade when encountering unseen or extreme poses.
-In future work, we plan to investigate more flexible garment representations and explore pose-independent anchor structures to better capture complex clothing dynamics and improve generalization to diverse motions.
-Figure 13: More animation results of avatars created with 8-image inputs. Reference image is one of the input images.
+Appendix E Limitations and Future Work
+A primary limitation of our approach is its dependence on the SMPL-X template mesh to initialize geometric tokens.
+While this provides a strong structural prior, it can limit reconstruction fidelity for subjects wearing loose or non-body-conforming garments, such as dresses and skirts.
+As shown in Figure 14, this manifests in failure cases where the model struggles to animate avatars wearing skirts during large lower-body movements.
+Furthermore, due to the limited diversity of large-motion poses in our training datasets, the model’s performance may decline when faced with unseen or extreme poses, which is also reflected in the failure cases presented.
+Additionally, since our model is primarily designed for sparse view input, we find that as the number of input views exceeds 16, the improvement in performance becomes marginal.
+In future work, we plan to investigate more flexible garment representations and explore pose-independent anchor structures to better capture complex clothing dynamics, enhancing generalization to diverse motions.
+Moreover, our aim is to design a framework that is more suitable for dense view inputs.
+Inputs Animation Results Inputs Animation Results
+Figure 14: Failure Cases. our model struggles to animate avatars wearing skirts during large-amplitude lower‑body movements, and has difficulty handling inputs with extreme or challenging poses.
+Reference Reconstruction Reference Reconstruction
+Figure 15: Human reconstruction in canonical space from 8 input images. The reference image is one of the inputs.
+Figure 16: More animation results of avatars created with 8-image inputs (Part I). Reference image is one of the input images.
+Figure 17: More animation results of avatars created with 8-image inputs (Part II). Reference image is one of the input images.